### v2

# Moving Least Squares without Quasi-Uniformity: A Stochastic Approach

Abstract
Local Polynomial Regression (LPR) and Moving Least Squares (MLS) are closely related nonparametric estimation methods, developed independently in statistics and approximation theory. While statistical LPR analysis focuses on overcoming sampling noise under probabilistic assumptions, the deterministic MLS theory studies smoothness properties and convergence rates with respect to the fill-distance (a resolution parameter). Despite this similarity, the deterministic assumptions underlying MLS fail to hold under random sampling. We begin by quantifying the probabilistic behavior of the fill-distance $h_{n}$ and separation $\delta_{n}$ of an i.i.d. random sample. That is, for a distribution satisfying a mild regularity condition, $h_{n}\propto n^{-1/d}\log^{1/d}(n)$ and $\delta_{n}\propto n^{-2/d}$. We then prove that, for MLS of degree k1k\!-\!1, the approximation error associated with a differential operator $Q$ of order $|m|\leq k-1$ decays as $h_{n}^{\,k-|m|}$ up to logarithmic factors, establishing stochastic analogues of the classical MLS estimates. Additionally, We show that the MLS approximant is smooth with high probability. Finally, we apply the stochastic MLS theory to manifold estimation. Assuming that the sampled Manifold is $k$-times smooth, we show that the Hausdorff distance between the true manifold and its MLS reconstruction decays as $h_{n}^{k}$, extending the deterministic Manifold-MLS guarantees to random samples. This work provides the first unified stochastic analysis of MLS, demonstrating that—despite the failure of deterministic sampling assumptions—the classical convergence and smoothness properties persist under natural probabilistic models.
1 Introduction
1.1 Background
Local Polynomial Regression (LPR) and Moving Least Squares (MLS) are two closely related nonparametric methods, both relying on weighted local polynomial fitting. While they emerged independently—LPR from statistics and MLS from numerical analysis—their mathematical formulations are essentially equivalent. This parallel development has led to two complementary traditions: one focused on statistical inference under random sampling, the other on numerical approximation under deterministic sampling.
In statistics, LPR was introduced as a framework for estimating functions and their derivatives from noisy observations. The method constructs a local polynomial approximation around a target point, with nearby data weighted more heavily than distant ones. Its development has been driven by classical statistical concerns such as asymptotic consistency, the bias–variance trade-off, and optimality in stochastic settings. Early work by Stone Stone (1982) established the global optimal rates of convergence for nonparametric estimators, showing how the smoothness of the underlying function and the dimension of the data determine achievable accuracy. Subsequent contributions, including those of Ruppert and Wand Ruppert and Wand (1994), provided refined asymptotic analyses of bias and variance, emphasizing the role of bandwidth selection and boundary effects. Within this framework, LPR was first studied for point estimation and later extended to interval estimation, consolidating its role as a central tool in nonparametric statistics.
In parallel, MLS was developed in computational mathematics as a meshfree method for function approximation Levin (1998). Like LPR, it employs weighted polynomial fitting, but its construction yields smooth shape functions with local support that enable continuous reconstruction across a domain. This property makes MLS especially suitable for geometric applications, such as surface reconstruction Levin (2004), solving partial differential equations Mirzaei (2015), and manifold approximation Sober and Levin (2020). The deterministic analyses of MLS rely on geometric sampling conditions, notably the notions of fill distance (sometimes referred to as the covering radius), the largest gap between sample points and the domain, and separation, the minimal distance between two distinct points. In this setting, the two quantities are comparable up to a constant factor, which provides a convenient framework for proving convergence and stability. Within this setting, Levin Levin (1998) demonstrated convergence rates for smooth functions, while Mirzaei Mirzaei (2015) extended the theory to derivatives. In particular, Mirzaei showed that if one applies a differential operator of order $m$ to both the MLS approximation and the true function, the difference between them decays like the fill distance raised to the power kmk-m, where $k$ is the smoothness order of the function. These results underscore the strength of MLS in preserving smoothness and geometric structure under deterministic sampling assumptions. More recently, Krieg and Sonnleitner Krieg and Sonnleitner (2024) relaxed the reliance on fill distance by showing that random points on convex domains are asymptotically optimal for Sobolev approximations. This suggests that the large gaps characteristic of stochastic sampling do not prevent optimal average convergence rates.
Although LPR and MLS arose from distinct motivations—statistical inference and numerical approximation, respectively—their equivalence suggests a unified perspective. On the one hand, the global approximation framework and geometric intuition of MLS offer a natural way to extend LPR to data supported on manifolds. On the other hand, the stochastic theory of LPR provides the probabilistic tools needed to analyze MLS when data are generated randomly. Yet a fundamental obstacle arises: the deterministic assumptions underlying MLS, such as fill distance and separation, do not hold under stochastic sampling. This prevents a direct transfer of results from one framework to the other.
The main contribution of this work is to bridge this gap. The deterministic theory of MLS relies on a geometric relation between the fill distance and the separation, but these assumptions are not satisfied when the data are drawn stochastically. Within this stochastic framework, we reestablish the central approximation results of Mirzaei, thereby extending his deterministic convergence results to random samples. In particular, we show that the rate of convergence for the error between derivatives of the MLS approximation and derivatives of the true function coincides with the optimal rate established by Stone Stone (1982) in statistics. Building on this, we further extend the theory of manifold approximation by MLS—originally formulated under deterministic assumptions—to the stochastic regime, demonstrating that the same convergence rates persist.
2 Preliminaries
We first recall some standard geometric concepts from the deterministic MLS framework.
Definition 1.
Let $\Omega$ be a domain in $\mathbb{R}^{d}$ , and consider sets of data points in $\Omega$. We say that $X=\left\{x_{i}\right\}_{i=1}^{I}$ is a $h-\delta$ set if:
  1. 1.
    $h$ is the ”fill distance” with respect to the domain $\Omega$,
    $h=\sup_{x\in\Omega}\min_{x_{i}\in X}\left\|x-x_{i}\right\|,$
  2. 2.
    The ”separation” is the largest $\delta$ such that
    $\left\|x_{i}-x_{j}\right\|\geq\delta,\quad\forall\,1\leq i<j\leq I.$
The next definition describes a condition on the sample points that ensures they are neither too close together nor too far apart, providing a foundation for all subsequent deterministic MLS results.
Definition 2.
(quasi-uniform sample set). A set of data sites $X=\left\{x_{1},\ldots,x_{n}\right\}$ is said to be quasi-uniform with respect to a domain $\Omega$ and a constant $c_{\mathrm{qu}}>0$ if
$\delta\leq h\leq c_{\mathrm{qu}}\delta$
where $h$ is the fill distance and $\delta$ is the separation.
While this definition serves as the foundation for the subsequent results in the deterministic setting, it does not hold in the stochastic setup, which required us to rework the proofs to adapt them to the stochastic conditions, as will be seen later.
We now recall some central results from the deterministic MLS framework, focusing on the construction of the local polynomial approximation and its convergence properties.
Let $\theta_{h}\in\mathcal{C}^{k}$ (continuously differentiable $k$ times over $\Omega$) be a weight function with compact support of size $s\cdot h$ of the shape
$\theta_{h}(x)=\Phi\left(\frac{x}{h}\right),$ (1)
where $\Phi:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a nonnegative function supported in the ball $B(0,s)$ (the ball centered at the origin with radius $s$), where $\Phi(x)>0$ for all $x\in B(0,s)$.
Let $\left\{x_{i}\right\}_{i=1}^{n}$ be a set of distinct scattered points in $\mathbb{R}^{d}$, and $\left\{f\left(x_{i}\right)\right\}_{i=1}^{n}$ be the corresponding sampled values of some function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Then, we define the local polynomial by $p^{*}_{\xi}$,
$p^{*}_{\xi}=\underset{p\in\Pi_{k-1}^{d}}{\operatorname{argmin}}\sum_{i=1}^{n}\left(p\left(x_{i}\right)-f\left(x_{i}\right)\right)^{2}\theta_{h}\left(\xi-x_{i}\right),$ (2)
where $\Pi_{k-1}^{d}$ is the space of polynomials of total degree k1k-1 in $\mathbb{R}^{d}$. The MLS function is defined by
$s^{\text{MLS}}_{f,X}(x):=p^{*}_{x}(0).$
Based on the above, Levin Levin (1998) showed that if $f\in\mathcal{C}^{k}(\Omega)$ then
$\|s^{\text{MLS}}_{f,X}(x)-f\|_{\Omega,\infty}<L\cdot h^{k},$
where $\|s^{\text{MLS}}_{f,X}(x)-f\|_{\Omega,\infty}=\sup_{x\in\Omega}|s^{\text{MLS}}_{f,X}(x)-f(x)|$, and $L$ is a constant. Another significant theoretical result was established in Mirzaei’s work Mirzaei (2015), which demonstrated that for some constant $C$ independent of $f$ and $h$, if $f$ is $k$-times differentiable, then
$\left|\partial^{\alpha}s^{\text{MLS}}_{f,X}(\hat{x})-\partial^{\alpha}f(\hat{x})\right|\leq C|f|_{\mathcal{C}^{k}(\hat{x})}h^{k-|\alpha|},$
where $\alpha=\left(\alpha_{1},\ldots,\alpha_{d}\right)$ is a multi-index, meaning an ordered tuple of nonnegative integers used to denote partial derivatives: $\partial^{\alpha}f=\frac{\partial^{|\alpha|}f}{\partial x_{1}^{\alpha_{1}}\ldots\partial x_{d}^{\alpha_{d}}}\text{, where }|\alpha|=\alpha_{1}+\cdots+\alpha_{d}.$
More generally, let $Q$ be a linear differential operator with constant coefficients, given by
$Q=\sum_{|\alpha|\leq k}q_{\alpha}\partial^{\alpha},$
where $q_{\alpha}$ are real constants for $|\alpha|\leq k$. Here, $\alpha=\left(\alpha_{1},\ldots,\alpha_{d}\right)$ denotes a multi-index of nonnegative integers. Its order is $|\alpha|=\alpha_{1}+\cdots+\alpha_{d}$, and for a smooth function $f$ we write $\partial^{\alpha}f=\frac{\partial^{|\alpha|}f}{\partial x_{1}^{\alpha_{1}}\cdots\partial x_{d}^{\alpha_{d}}}$. Let $m$ denote the order of $Q$, defined by
$m=\max\left(|\alpha|:|\alpha|\leq k-1\text{ and }q_{\alpha}\neq 0\right).$
(For example, if $Qf=f$, then $m=0$; while if $Qf=\partial^{2}f/\partial x_{1}^{2}+\cdots+\partial^{2}f/\partial x_{1}\partial x_{2}$, then $m=2$).
With these definitions and results in place, we are now ready to extend the MLS framework from the deterministic setting to stochastic samples, which will be the focus of the next section.
3 Main Results
Theorem 1.
For a distribution $\mathbb{P}$ over $\Omega$ that satisfies Definition 4, let $X_{n}=\{x_{1},x_{2},...,x_{n}\}$ be a set of $n$ iidiid samples of $\mathbb{P}$. Denote by $h_{n}$ the fill distance of $X_{n}$ (with respect to $\Omega$), and by $\delta_{n}$ the separation of $X_{n}$. Then,
$h_{n}=O_{P}(n^{-\frac{1}{d}}\ln^{\frac{1}{d}}(n))\quad\mbox{and}\quad\delta_{n}=\Omega_{P}(n^{-\frac{2}{d}}).$ (3)
Proof.
The proof follows immediately from Lemma 7 and Lemma 15 below. ∎
As a result, we needed to adapt the proofs of Mirzaei Mirzaei (2015) so that they hold under the stochastic conditions of Theorem 3. Ultimately, we are able to establish the following theorem, which serves as the stochastic analogue of Mirzaei’s result.
Theorem 2.
Let $X_{n}$ be a sample set of size $n$ drawn from a nicely behaving distribution (Definition 4) on $\Omega$, which satisfies the interior cone condition (Definition 3). Let $F_{n}$ be the corresponding sampled values of some function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and let $s^{\text{MLS}}_{F_{n},X_{n}}(x)$ be the value of the (k1)(k-1) degree MLS function at the point $x$, then for $\hat{x}\in\Omega$
$P\!\left(\left|Qs^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-Qf(\hat{x})\right|\geq\mathcal{K}\,h_{n}^{k-|m|}\log n\right)\leq\frac{2}{n^{c}}+\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}+\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}$ (4)
i.e.,
$\lim_{n\rightarrow\infty}P\left(\left|Qs^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-Qf(\hat{x})\right|\geq\mathcal{K}\cdot h_{n}^{k-|m|}\cdot(\log(n))\right)=0,$ (5)
where $\mathcal{K}$ is a positive constant depends on $\varepsilon$, $Q$ is a linear differential operator of degree $\leq k$, $|m|\leq k-1$ is some specific multi-index, and $c,c^{\prime},c_{\mathcal{Q}}$ are some positive constants.
See proof in Section 5.
Theorem 3.
Given $\theta_{h}\in\mathcal{C}^{k}$, the MLS approximation $s^{\text{MLS}}_{F_{n},X_{n}}(x)\in\mathcal{C}^{k}(\Omega)$ with probability at least $1-\frac{2}{n^{\tilde{c}_{\lambda}}}$, where $\tilde{c}_{\lambda}$ is a positive constant.
See proof in Section 6.
Remark 1.
The following statement is presented here in its final form, although some of the definitions and notations it involves will only be introduced later (see Section 7).
Theorem 4.
Let $\mathcal{M}^{h_{n}}$ denote the Manifold-MLS approximation of degree k1k-1. Assume that $\mathcal{M}^{h_{n}}\in\mathcal{C}^{k}$. Then, there exists a constant $s>0$, independent of $h_{n}$, such that the Manifold-MLS approximation using a radial weight function with finite support of size $sh_{n}$ satisfies, for sufficiently small $h_{n}$, the error bound
$\left\|\mathcal{M}^{h_{n}}-\mathcal{M}\right\|_{\text{Hausdorff}}=O_{P}(\log nh_{n}^{k}).$
See proof in Section 7.
4 Bridging the gap between deterministic and stochastic MLS
This section lays the probabilistic foundations required to rigorously compare deterministic and stochastic MLS frameworks. We begin by introducing two central definitions: the interior cone condition, which imposes a geometric regularity on the domain, and a family of distributions with well-behaved sampling properties. Under these assumptions, we establish that the convergence rate of the fill distance differs from that of the separation distance—a key distinction from the deterministic setting. We then derive properties of how the number of sample points in local neighborhoods grows with $n$, and other structural properties of the random point cloud, such as local point density and inter-point distances, that are essential for the analysis.
Definition 3.
A set $\Omega\subseteq\mathbb{R}^{d}$ is said to satisfy an interior cone condition if there exists an angle $\theta\in(0,\pi/2)$ and a radius $r>0$ such that for every $x\in\Omega$ a unit vector $\xi(x)$ exists such that the cone
$C(x,\xi(x),\theta,r):=\left\{x+\lambda y:y\in\mathbb{R}^{d},\|y\|_{2}=1,y^{T}\xi(x)\geq\cos\theta,\lambda\in[0,r]\right\}$ (6)
is contained in $\Omega$.
Remark 2.
If the domain has no boundary, then the interior cone condition is automatically satisfied, and all proofs in the article remain valid . However, if the domain has a boundary, the interior cone condition becomes necessary. The interior cone condition (see Definition 3) ensures the existence of an angle $\theta$ and a radius $r$ such that a cone of positive volume can be positioned at any point in the domain $\Omega$. As $n\to\infty$, the number of data points contained within such a cone also diverges to infinity. In particular, the number of samples in the cone is proportional to $n$, with the proportionality factor given by the angular ratio $\theta/(2\pi)$. This guarantees that no substantial difficulty arises near the boundary.
Having introduced the geometric assumptions on the domain, we now formalize the probabilistic setting. Let $(\Omega,p(x))$ be a probability space. Assume $\Omega$ is a closed compact and satisfies an interior cone condition. Let $p(x)$ be the probability density function. We assume $p(x)>0$ $\forall x\in\Omega$.
Definition 4.
(”nicely behaving distribution”). Let $\mu$ denote the uniform measure over $\Omega$, induced by the Lebesgue measure. A probability density function $p(x)$ on $\Omega$ is called nicely behaving distribution, if there exist constants $0<\underline{c}\leq\bar{c}<\infty$ such that $\underline{c}\cdot\mu\leq p(x)\leq\bar{c}\cdot\mu$.
We now turn to the proof of the Theorem 3 stated above. The proof is divided into two lemmas: the first establishes the asymptotic behavior of the fill distance, while the second addresses the asymptotic behavior of the separation. Together, these results complete the proof of the theorem and highlight the fundamental distinction between the stochastic and deterministic settings. Specifically, while the deterministic framework typically assumes quasi-uniformity (see Definition 2), this property fails in the stochastic case.
Lemma 1.
For a distribution $\mathbb{P}$ over $\Omega$ that satisfies Definition 4, let $X_{n}=\{x_{1},x_{2},...,x_{n}\}$ be a set of $n$ iidiid samples of $\mathbb{P}$. Denote by $h_{n}$ the fill distance with respect to $X_{n}$. Then,
$h_{n}=O_{P}(n^{-\frac{1}{d}}\ln^{\frac{1}{d}}(n)).$ (7)
Remark 3.
A similar upper bound for $h_{n}$ was first established in Reznikov and Saff (2016). We include this proof for completeness.
Proof.
From the intrior cone condition we obtain that $\text{Vol}_{d}(B(x,\varepsilon)\cap\Omega)>0$ for all $x\in\Omega$ and all real $\varepsilon>0$, where $B(x,\varepsilon)$ is the open ball in $\mathbb{R}^{d}$ of radius $\varepsilon$ centered at $x$. Then the condition $p(x)>0$ for all $x\in\Omega$ implies that $P(x_{1}\in B(x,\varepsilon)\cap\Omega)>0$ and hence $P(x_{1}\notin B(x,\varepsilon)\cap\Omega)<1$, for all $x\in\Omega$ and all real $\varepsilon>0$. In addition, by the compactness of $\Omega$, for any real $\varepsilon>0$ there is a finite minimal covering set $F_{\varepsilon}\subseteq\Omega$ such that $\Omega\subseteq\bigcup_{x\in F_{\varepsilon}}B(x,\varepsilon)$. The cardinality of $F_{\varepsilon}$ represents the smallest number of closed balls with centers in $\Omega$ and radii $\varepsilon$ whose union covers $\Omega$. We denote the covering number by $\mathcal{N}_{\varepsilon}=|F_{\varepsilon}|$. Then, for each real $\varepsilon>0$,
$P\left(h_{n}\geq 2\varepsilon\right)\leq P\left(\bigcup_{x\in F_{\varepsilon}}\bigcap_{i=1}^{n}\left\{x_{i}\notin B(x,\varepsilon)\right\}\right)\leq\sum_{x\in F_{\varepsilon}}P\left(x_{1}\notin B(x,\varepsilon)\right)^{n}.$ (8)
Note that as $n\rightarrow\infty$ this propbabilty goes to zero, hence, $h_{n}\rightarrow 0$ in probability.
Now, in order to find the probabilistic rate of decay of $h_{n}$, we first bound the covering number of the domain. It is shown in Proposition 4.2.12 of Vershynin (2018) that for any set $K\subset\mathbb{R}^{d}$ and any $\varepsilon>0$, the covering number $\mathcal{N}_{\varepsilon}$ satisfies
$\frac{\operatorname{vol}_{d}(K)}{\operatorname{vol}_{d}(B(x,\varepsilon))}\leq\mathcal{N}_{\varepsilon}\leq\frac{\operatorname{vol}_{d}(K+B(x,\varepsilon))}{\operatorname{vol}_{d}(B(x,\varepsilon))},$ (9)
where $\operatorname{vol}_{d}(\cdot)$ denotes $d$-dimensional Lebesgue measure and ++ denotes the Minkowski sum. Now, if we denote by $P(B(x,\varepsilon))$ the probability to have a sample in $B(x,\varepsilon)$, then according to the assumptions $\int_{B(x,\varepsilon)}\underline{c}d\mu\leq\int_{B(x,\varepsilon)}dp\leq\int_{B(x,\varepsilon)}\bar{c}d\mu$, it follows that $P(B(x,\varepsilon))\geq\underline{c}\cdot P_{U}(B(x,\varepsilon))$, where $P_{U}$ is the probability density function (pdf) of the uniform distribution over $\Omega$. Thus, the probability that there are no samples in $B(x,\varepsilon)$, is less or equal $1-\underline{c}\cdot P_{U}(B(x,\varepsilon))$, and therefore we get that
$\sum_{x\in F_{\varepsilon}}P\left(x_{1}\notin B(x,\varepsilon)\right)^{n}\leq\sum_{i=1}^{\mathcal{N}_{\varepsilon}}\left(1-\underline{c}\cdot P_{U}(B(x,\varepsilon))\right)^{n}=\mathcal{N}_{\varepsilon}\left(1-\underline{c}\cdot P_{U}(B(x,\varepsilon))\right)^{n}.$ (10)
(Note that $h_{n}\rightarrow 0$ as $n\rightarrow\infty$). Since $\Omega$ is compact, its volume exists and is finite. In addition, for a very small $\varepsilon$, the contribution of $B(x,\varepsilon)$ to the total volume is marginal, and we denote this total volume by $vol_{d}(\Omega+B(x,\varepsilon))=\kappa$. Using this notation with substituting Equation (9), we get that
$P\left(h_{n}\geq 2\varepsilon\right)\leq\frac{\kappa}{\varepsilon^{d}}\left(1-\underline{c}\cdot P_{U}(B(x,\varepsilon))\right)^{n}.$ (11)
In particular, we have that $P_{U}(B(x,\varepsilon))=C\cdot\varepsilon^{d}$, where $C=\frac{\pi^{\frac{d}{2}}}{\text{Vol}_{d}(\Omega)\Gamma(\frac{d}{2}+1)}$ and $\Gamma$ is Euler’s Gamma function. Hence, if $c=\underline{c}\cdot C$, then
$P\left(h_{n}\geq 2\varepsilon\right)\leq\frac{\kappa}{\varepsilon^{d}}\left(1-c\varepsilon^{d}\right)^{n}.$ (12)
Now, in order to find convergence rate of $h_{n}$, we need to express $\varepsilon$ in terms of n. We know that $1-x\leq e^{-x}$ for all $x$. So, if we let $x=c\varepsilon^{d}$, then
$1-c\varepsilon^{d}\leq e^{-c\varepsilon^{d}}.$
Using this inequality in Equation (8), we find that
$\frac{\kappa}{\varepsilon^{d}}\cdot\left(1-c\varepsilon^{d}\right)^{n}\leq\frac{\kappa}{\varepsilon^{d}}\cdot e^{-nc\varepsilon^{d}}.$
Furthermore, applying $e^{-y}\leq y\cdot e^{-y}$ for $y\geq 1$, then for $y=nc\varepsilon^{d}$ we find that
$\frac{\kappa}{\varepsilon^{d}}\left(e^{-nc\varepsilon^{d}}\right)\leq\frac{\kappa}{\varepsilon^{d}}\cdot\left(nc\varepsilon^{d}\cdot e^{-nc\varepsilon^{d}}\right).$
So now we require that $\frac{\kappa}{\varepsilon^{d}}\cdot\left(nc\varepsilon^{d}\cdot e^{-nc\varepsilon^{d}}\right)\leq\rho_{h}$ (where $0<\rho_{h}<1$ since it represents a probability), and if we solve the later inequality we find that
$\varepsilon\geq\frac{\ln^{\frac{1}{d}}(n\cdot\frac{\kappa c}{\rho_{h}})}{n^{\frac{1}{d}}\cdot c^{\frac{1}{d}}}.$ (13)
For any $d\geq 1$, exists $K_{h}$, a real positive constant, and $N_{h}$ such that for any $n\geq N_{h}$,
$\varepsilon\geq K_{h}\cdot n^{-\frac{1}{d}}\cdot\ln^{\frac{1}{d}}(n).$ (14)
We can then rewrite Equation (12) as
$P(h_{n}\geq 2\varepsilon)\leq\rho_{h},$
which is equivalent to
$P\left(\left|\frac{h_{n}}{\varepsilon}\right|>2\right)\leq\rho_{h}.$
Substituting the lower bound of $\varepsilon$ from Equation (14), keeps this inequality valid. Hence, exists a finite $N_{h}>0$ such that
$P\left(\left|\frac{h_{n}}{K_{h}\cdot n^{-\frac{1}{d}}\cdot\ln^{\frac{1}{d}}(n)}\right|>2\right)\leq\rho_{h},\,\,\forall n>N_{h},$
or, equivalently, $h_{n}=O_{P}(n^{-\frac{1}{d}}\ln^{\frac{1}{d}}(n))$. ∎
Lemma 2.
For a distribution $\mathbb{P}$ over $\Omega$ that satisfies Definition 4, let $X_{n}=\{x_{1},x_{2},...,x_{n}\}$ be a set of $n$ iidiid samples of $\mathbb{P}$. Denote by $\delta_{n}$ the separation of $X_{n}$. Then,
$\delta_{n}=\Omega_{P}(n^{-\frac{2}{d}}).$ (15)
Remark 4.
A version of the bound in Lemma 15 was previously established by Brauchart et al. (2018) for the specific case of a sphere under a uniform distribution. Our result generalizes this bound to any compact domain in $\mathbb{R}^{n}$ and relaxes the distribution requirement to almost uniform distributions.
Proof.
Recall that $P(B(x,\varepsilon))$ is the probability to have an $\varepsilon$-ball centered at $x$ with a sample in it. We’ve shown that this probability is bounded from above by $\bar{c}\cdot P_{U}(B(x,\varepsilon))$, and bounded from below by $\underline{c}\cdot P_{U}(B(x,\varepsilon))$. Note that $P(\delta_{n}>\varepsilon)$ represents the probability that no two samples from $X$ occupy the same $\varepsilon$-ball around either of them. Thus, considering $\varepsilon$-balls around all samples and using the fact that these balls may not be disjoint, we get the following bound
$P(\delta_{n}>\varepsilon)\geq\prod_{i=1}^{n-1}(1-\sum_{j=1}^{i}P(B(x_{j},\varepsilon)).$ (16)
Hence,
$P(\delta_{n}<\varepsilon)\leq 1-\prod_{i=1}^{n-1}(1-\sum_{j=1}^{i}P(B(x_{j},\varepsilon)).$ (17)
Using the upper bound of $P(B(x,\varepsilon))$, we obtain that
$1-\prod_{i=1}^{n-1}(1-\sum_{j=1}^{i}P(B(x_{j},\varepsilon))\leq 1-\prod_{i=1}^{n-1}(1-i\cdot\bar{c}\cdot P_{U}(B(x_{i},\varepsilon))),$ (18)
if $c^{\prime}=\bar{c}\cdot C$, then by previous results,
$1-\prod_{i=1}^{n-1}(1-i\cdot P\left(B(x_{i},\varepsilon)\right))\leq 1-\prod_{i=1}^{n-1}(1-i\cdot c^{\prime}\varepsilon^{d}).$ (19)
Hence,
$P\left(\delta_{n}<\varepsilon\right)\leq 1-\prod_{i=1}^{n-1}(1-i\cdot c^{\prime}\varepsilon^{d}).$ (20)
Now,
$1-\prod_{i=1}^{n-1}\left(1-i\cdot c^{\prime}\varepsilon^{d}\right)\leq 1-\prod_{i=1}^{n-1}\left(1-nc^{\prime}\varepsilon^{d}\right)=1-\left(1-nc^{\prime}\varepsilon^{d}\right)^{n},$ (21)
i.e.,
$P\left(\delta_{n}<\varepsilon\right)\leq 1-\left(1-nc^{\prime}\varepsilon^{d}\right)^{n}.$ (22)
Fix any $\rho_{\delta}\in(0,1)$. We seek to find a sequence $\varepsilon(n)$ such that the right-hand side of (22) is bounded by $\rho_{\delta}$. Let $\varepsilon(n)$ be defined by setting:
$1-(1-nc^{\prime}\varepsilon^{d})^{n}=\rho_{\delta}.$ (23)
Solving for $\varepsilon(n)$ yields:
$\varepsilon(n)=\left(\frac{1-(1-\rho_{\delta})^{1/n}}{nc^{\prime}}\right)^{1/d}.$ (24)
To obtain a clean bound for (24), we observe that $(1-\rho_{\delta})^{1/n}=\exp\left(\frac{1}{n}\log(1-\rho_{\delta})\right)$. Applying the inequality $e^{-x}\geq 1-x$ for $x\geq 0$, where $x=\frac{1}{n}\log\left(\frac{1}{1-\rho_{\delta}}\right)$, we have:
$(1-\rho_{\delta})^{1/n}\geq 1-\frac{\log(1/(1-\rho_{\delta}))}{n}.$ (25)
Rearranging this inequality gives:
$1-(1-\rho_{\delta})^{1/n}\leq\frac{\log(1/(1-\rho_{\delta}))}{n}.$ (26)
Substituting this into (24), we define the constant $K_{\delta}$ as:
$K_{\delta}:=\left(\frac{\log(1/(1-\rho_{\delta}))}{c^{\prime}}\right)^{1/d}.$ (27)
It follows that for all $n\geq 1$:
$\varepsilon(n)\leq K_{\delta}\,n^{-2/d}.$ (28)
Consequently, substituting this bound back into (22), we obtain:
$P\left(\delta_{n}<K_{\delta}\,n^{-2/d}\right)\leq\rho_{\delta},\quad\forall n\geq 1.$ (29)
By the definition of order in probability for lower bounds, this strictly implies:
$\delta_{n}=\Omega_{P}(n^{-2/d}).$ (30)
Corollary 1.
Since $h_{n}=O_{P}(n^{-\frac{1}{d}}\ln^{\frac{1}{d}}(n))$ and $\delta_{n}=\Omega_{P}(n^{-\frac{2}{d}})$ we get that the quasi-uniform property does not hold (see Definition 2).
Corollary 1 poses a challenge, as the quasi-uniform property is a common assumption, particularly in the deterministic MLS method. Consequently, the proofs by Levin Levin (1998), Davoud Mirzaei (2015), and Wendland Wendland (2005) cannot be directly applied to a stochastically generated data set, necessitating modifications to their arguments. Our proposed solution is to prove Theorem 2, assuming that the stochastic data is well behaved distribution, as defined in Definition 4.
As a preliminary step, we first establish a fundamental lemma that captures a key asymptotic property of the sample size as $n\rightarrow\infty$, which underpins each of these results.
Lemma 3.
Let $D_{n}$ be a ball of radius $R_{n}=C_{D}\cdot\left(\frac{\log n}{n}\right)^{1/d}$, centered at some point $x\in\Omega$ and let
$N_{D_{n}}=\sum_{i=1}^{n}\mathbf{1}_{\{x_{i}\in D_{n}\}}$
denote the number of sample points lying in $D_{n}$. Then, there exist finite constants $0<\gamma_{1}\leq\gamma_{2}$ such that for some constant $c>0$,
$P\left(\gamma_{1}\log n\ \leq\ N_{D_{n}}\ \leq\ \gamma_{2}\log n\right)\geq 1-\frac{2}{n^{c}},$
that is,
$\lim_{n\to\infty}P\left(\gamma_{1}\log n\ \leq\ N_{D_{n}}\ \leq\ \gamma_{2}\log n\right)=1.$
Proof.
Let $N_{D_{n}}=\sum_{i=1}^{n}\mathbf{1}_{\{x_{i}\in D_{n}\}}$. Each indicator variable $\mathbf{1}_{\{x_{i}\in D_{n}\}}$ is Bernoulli with success probability $p_{n}=P(x_{i}\in D_{n})$. Then $N_{D_{n}}\sim\text{Bin}(n,p_{n})$, and:
$\mathbb{E}N_{D_{n}}=np_{n}=n\int_{D_{n}}p(x)dx.$
Since $D_{n}$ is a ball of radius $R_{n}=C_{D}\cdot\left(\frac{\log n}{n}\right)^{1/d}$, its volume is
$\operatorname{Vol}(D_{n})\propto R_{n}^{d}\propto\frac{\log n}{n}.$
Because $p(x)$ is bounded below by a constant $\underline{c}>0$ on its support, we obtain:
$\mathbb{E}N_{D_{n}}\geq n\cdot\int_{D_{n}}\underline{c}\,dx=\underline{c}\cdot n\cdot\operatorname{Vol}(D_{n})\geq c_{1}\cdot\log n$ (31)
for some constant $c_{1}>0$. Likewise, since $p(x)\leq\bar{c}$, we also get:
$\mathbb{E}N_{D_{n}}\leq\bar{c}\cdot n\cdot\operatorname{Vol}(D_{n})\leq c_{2}\cdot\log n.$ (32)
We now apply the Chernoff bound: If $X\sim\text{Bin}(n,p)$ with $\mu=\mathbb{E}X$, then for any $0<\delta<1$,
$P\left(|X-\mu|\geq\delta\mu\right)\leq 2\exp\left(-\frac{\delta^{2}\mu}{3}\right).$
Applying this to $N_{D_{n}}$ with $\mu=\mathbb{E}N_{D_{n}}$, we obtain:
$P\left(\left|N_{D_{n}}-\mathbb{E}N_{D_{n}}\right|\geq\delta\mathbb{E}N_{D_{n}}\right)\leq 2\exp\left(-c\log n\right)=\frac{2}{n^{c}},$
for some constant $c>0$. Thus,
$P\left((1-\delta)\mathbb{E}N_{D_{n}}\leq N_{D_{n}}\leq(1+\delta)\mathbb{E}N_{D_{n}}\right)\geq 1-\frac{2}{n^{c}}.$ (33)
Denote $\gamma_{1}=(1-\delta)c_{1}$ and $\gamma_{2}=(1+\delta)c_{2}$, and use the inequalities in (31),(32) we obtain that
$P\left(\gamma_{1}\log n\ \leq\ N_{D_{n}}\ \leq\ \gamma_{2}\log n\right)\geq 1-\frac{2}{n^{c}},$ (34)
which implies that
$\lim_{n\to\infty}P\left(\gamma_{1}\log n\ \leq\ N_{D_{n}}\ \leq\ \gamma_{2}\log n\right)=1.$
We denote by $\mathscr{E}_{I}$ the event
$\mathscr{E}_{I}:=\{\#I_{B(\hat{x},sh_{n})}\in[\gamma_{1}\log n,\gamma_{2}\log n]\}.$ (35)
We wish to emphasize an important point. As the total number of samples increases, the number of samples contained in any fixed ball tends to infinity. Consequently, the quasi-uniformity property fails. This is precisely the obstruction that prevents a direct application of the techniques developed in Levin (1998), Mirzaei (2015), Wendland (2005), and Sober and Levin (2020). In the stochastic setting, however, this phenomenon is not only unavoidable but also advantageous, since it is natural for the quality of the approximation to improve as the number of samples grows. By adapting the proof techniques to accommodate this behavior, together with our results, the difficulty is resolved, and the conclusions obtained in the deterministic framework can be retained.
5 Proof of Theorem 2
We first establish Lemma 4, which is a refined version of Theorem 2. In this lemma, the analysis is carried out for individual partial derivatives of fixed order, represented by multi-index derivatives. This allows us to treat each derivative separately and obtain precise bounds. We then derive Theorem 2 as a direct consequence by applying the lemma to general linear differential operators with constant coefficients.
Lemma 4.
Let $X_{n}$ be a sample set of size $n$ drawn from a nicely behaving distribution (Definition 4) on $\Omega$, which satisfies the interior cone condition (Definition 3). Let $F_{n}$ be the corresponding sampled values of some function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$, and let $s^{\text{MLS}}_{F_{n},X_{n}}(x)$ be the value of the (k1)(k-1) degree MLS function at a point $x$, then for any $\hat{x}\in\Omega$
$P\left(\left|\partial^{\alpha_{0}}s^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-\partial^{\alpha_{0}}f(\hat{x})\right|\leq\mathcal{K}\cdot h_{n}^{k-|\alpha_{0}|}\cdot\log(n)\right)\geq 1-\frac{2}{n^{c}}-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$ (36)
i.e.,
$\lim_{n\rightarrow\infty}P\left(\left|\partial^{\alpha_{0}}s^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-\partial^{\alpha_{0}}f(\hat{x})\right|\leq\mathcal{K}\cdot h_{n}^{k-|\alpha_{0}|}\cdot\log(n)\right)=1,$ (37)
where $\mathcal{K}$ is a positive constant depends on $\varepsilon$, $|\alpha_{0}|\leq k-1$ is some multi-index, and $c,c^{\prime},c_{\mathcal{Q}}$ are some positive constants.
Proof.
Given a set of scattered data points $X_{n}=\left\{x_{i}\right\}_{i=1}^{n}\subset\mathbb{R}^{d}$, we define the neighborhood index set around a fixed but arbitrary point $\hat{x}\in\Omega$:
$I_{B\left(\hat{x},sh_{n}\right)}=\left\{i:x_{i}\in B\left(\hat{x},sh_{n}\right)\right\},$
where $B\left(\hat{x},sh_{n}\right)=\left\{x\in\mathbb{R}^{d}:\|x-\hat{x}\|\leq sh_{n}\right\}$. In the following, we work under the assumption that event $\mathscr{E}_{I}$ hold, i.e., by (34), the expected number of samples in the index set $I_{B(\hat{x},sh_{n})}$ is contained in $\gamma_{1}\log n,\gamma_{2}\log n]$ with probability at least $1-\frac{2}{n^{c}}$.
In Levin (1998)it was shown that the MLS approximation is given by
$s^{\text{MLS}}_{F_{n},X_{n}}(x)=\sum_{i\in I_{B\left(\hat{x},sh_{n}\right)}}a_{i}^{*}(x)f\left(x_{i}\right),$
where the coefficients $a_{i}^{*}(x)$ are called shape functions, and obtained by minimizing the quadratic form
$a_{i}^{*}(x)=\arg\min_{a_{i}(x)}\sum_{i\in I_{B\left(\hat{x},sh_{n}\right)}}a_{i}(x)^{2}\frac{1}{\theta_{h}\left(x_{i}-x\right)},$
under the constraints
$\sum_{i\in I_{B\left(\hat{x},sh_{n}\right)}}a_{i}(x)p\left(x_{i}\right)=p(x),\quad\forall p\in\Pi_{k-1}\left(\mathbb{R}^{d}\right),$
and $\theta_{h}(x_{i}-x)$ is a weight function satisfies the properties in (1). Let $A$ denote the number of multi-indices $\alpha=(\alpha_{1},\dots,\alpha_{d})$ with nonnegative entries such that $|\alpha|\leq k$. It is shown in Mirzaei (2015) that the shape functions can be expressed in the following form:
$a_{i}^{*}(x)=\theta_{h}(x_{i}-x)\sum_{|\alpha|\leq k-1}\eta_{\alpha}(x)\frac{\left(x_{i}-x\right)^{\alpha}}{h_{n}^{|\alpha|}},\quad i=1,2,\ldots,n.$ (38)
where $\eta_{\alpha}(x)$ are the components of $\eta(x)\in\mathbb{R}^{A}$, the solution of the positive definite system
$\mathscr{A}_{n}(x)\eta(x)=\boldsymbol{p}(x),$
$\boldsymbol{p}(x)\in\mathbb{R}^{A}$ is the vector whose components are $p_{\alpha}(x)$,
$p_{\alpha}(x)=\left\{\frac{(x-\hat{x})^{\alpha}}{h_{n}^{|\alpha|}}\right\}_{0\leq|\alpha|\leq k-1},$
and $\mathscr{A}_{n}(x)\in\mathbb{R}^{A\times A}$ is a matrix with entries $\mathscr{A}_{n\alpha\beta}(x)$ where,
$\mathscr{A}_{n\alpha\beta}(x)=\sum_{i=1}^{n}\theta_{h}(x_{i}-x)p_{\alpha}(x)\cdot p_{\beta}(x),\quad\alpha,\beta=1,\ldots,A,$
which equals,
$\mathscr{A}_{n\alpha\beta}(x)=\sum_{i=1}^{n}\theta_{h}(x_{i}-x)\frac{(x_{i}-x)^{\alpha}}{h_{n}^{|\alpha|}}\cdot\frac{(x_{i}-x)^{\beta}}{h_{n}^{|\beta|}},\quad\alpha,\beta=1,\ldots,A,$ (39)
with $\eta(x),\boldsymbol{p}(x)\in\mathbb{R}^{A}$. To illustrate the approximation of a function by the stochastic MLS operator, we follow Levin’s approach Levin (1998). Given an arbitrary $\hat{x}\in\Omega$, we define the MLS error as
[…]
By adding and subtracting a polynomial $p(x)$ expressed in the monomial basis scaled by $h_{n}$ and centered at $\hat{x}$, we get
[…]
Denote by $p_{f}$ the taylor polynomial of $f(x)$ around $\hat{x}$. We have:
$\max_{x\in B(\hat{x},sh_{n})}|f(x)-p_{f}(x)|\leq C|f|_{\mathcal{C}^{k}\left(B\left(\hat{x},sh_{n}\right)\right)}h_{n}^{k},$
where the seminorm, $|f|_{\mathcal{C}^{k}\left(B\left(\hat{x},sh_{n}\right)\right)}$, is defined by
$|f|_{\mathcal{C}^{k}\left(B\left(\hat{x},sh_{n}\right)\right)}\triangleq\max_{\begin{subarray}{c}|\alpha|=k\\ x\in B\left(\hat{x},sh_{n}\right)\end{subarray}}\left|\partial^{\alpha}f(x)\right|,$
with $\alpha=\left(\alpha_{1},\ldots,\alpha_{d}\right)$ denoting a multi-index of order $|\alpha|=k$, and
$\partial^{\alpha}f(x)=\partial_{x_{1}}^{\alpha_{1}}\partial_{x_{2}}^{\alpha_{2}}\cdots\partial_{x_{d}}^{\alpha_{d}}f(x).$
Using $p_{f}$ as the polynomial in Equation 40 we get
$\left|E_{F_{n},X_{n}}(\hat{x})\right|\leq C|f|_{\mathcal{C}^{k}\left(B\left(\hat{x},sh_{n}\right)\right)}h_{n}^{k}.$
Since the focus is on measuring the error in the derivatives, we obtain
$\partial^{\alpha}E_{F_{n},X_{n}}(\hat{x})=\sum_{i\in I_{B\left(\hat{x},h_{n}\right)}}\left(\partial^{\alpha}a_{i}^{*}(\hat{x})\right)\left(f\left(x_{i}\right)-p_{f}\left(x_{i}\right)\right)+\partial^{\alpha}(p_{f}(x)-f(x)).$
Note that in the current case, $\partial^{\alpha}$ is scalar-valued as $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Then, for a fixed but arbitrary $\hat{x}\in\Omega$ we get
$\left|\partial^{\alpha}E_{F_{n},X_{n}}(\hat{x})\right|\leq\max_{x\in B(\hat{x},sh_{n})}|f(x)-p_{f}(x)|\sum_{i\in I_{B\left(\hat{x},sh_{n}\right)}}\left|\partial^{\alpha}a_{i}^{*}(\hat{x})\right|+\left|\partial^{\alpha}p_{f}(x)-\partial^{\alpha}f(x)\right|.$ (41)
Since $p_{f}$ is the Taylor expansion of $f(x)$ at $\hat{x}$ and $\|\hat{x}-x\|\leq sh_{n}$ we get that
$\left|\partial^{\alpha}p_{f}(x)-\partial^{\alpha}f(x)\right|\leq c_{1}|f|_{\mathcal{C}^{k}(B(\hat{x},sh_{n}))}h_{n}^{k-|\alpha|}.$ (42)
Given the event $\mathscr{E}_{I}$ hold, we showed in Lemma 9 that for all $x\in B(\hat{x},sh_{n})$,
$P\left(\sum_{i\in I_{B\left(x,sh_{n}\right)}}\left|\partial^{\alpha}a_{i}^{*}(x)\right|\leq C_{\alpha}^{\prime}\cdot\log(n)\cdot h_{n}^{-|\alpha|}\Big|\mathscr{E}_{I}\right)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$ (43)
where $C^{\prime}_{\alpha}$ depends on $\varepsilon$. Hence the probability of the event in (43) intersects with $\mathscr{E}_{I}$ is bounded by,
$P\left(\sum_{i\in I_{B\left(x,sh_{n}\right)}}\left|\partial^{\alpha}a_{i}^{*}(x)\right|\leq C_{\alpha}^{\prime}\cdot\log(n)\cdot h_{n}^{-|\alpha|}\right)\geq 1-\frac{2}{n^{c}}-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}.$ (44)
Substituting the bounds from Equations (42) and (43) into Equation (41), and using the fact that $\max_{x\in B(\hat{x},h_{n})}|f(x)-p_{f}(x)|$ is bounded by $h_{n}^{k}$ up to a positive constant (Taylor remainder estimate), we evaluate at $\hat{x}$ for a specific $\alpha_{0}$, and we obtain that for some positive constant $\mathcal{K}$ depends on $\varepsilon$
$P\left(\left|\partial^{\alpha_{0}}s^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-\partial^{\alpha_{0}}f(\hat{x})\right|\leq\mathcal{K}\cdot h_{n}^{k-|\alpha_{0}|}\cdot\log(n)\right)\geq 1-\frac{2}{n^{c}}-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}.$
That is,
$\lim_{n\rightarrow\infty}P\left(\left|\partial^{\alpha_{0}}s^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-\partial^{\alpha_{0}}f(\hat{x})\right|\leq\mathcal{K}\cdot h_{n}^{k-|\alpha_{0}|}\cdot\log(n)\right)=1,$
as desired. ∎
Proof of Theorem 2.
Let $Q$ be a linear differential operator of order $m$, which can be written as
$Q=\sum_{|\alpha|\leq m}q_{\alpha}\,\partial^{\alpha},$
where the coefficients $q_{\alpha}$ are real constants. Let
$M_{Q}:=\{\alpha:\,|\alpha|\leq m,\ q_{\alpha}\neq 0\},$
which is a finite index set. Define the vector $\mathscr{Q}_{n}\in\mathbb{R}^{\#(M_{Q})}$ by
$(\mathscr{Q}_{n})_{\alpha}=\frac{\alpha!\,q_{\alpha}}{h_{n}^{|\alpha|}},\qquad\alpha\in M_{Q}.$
With this notation, the operator $Q$ can be expressed as
$Q=\mathscr{Q}_{n}^{\prime}\,\partial^{\alpha}.$
By Lemma 4, for each fixed $\alpha\in M_{Q}$,
$P\!\left(\left|\partial^{\alpha}s^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-\partial^{\alpha}f(\hat{x})\right|\geq\mathcal{K}\,h_{n}^{k-|\alpha|}\log n\right)\leq\frac{2}{n^{c}}+\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}+\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}.$
Applying the union bound over the finite set $M_{Q}$, we obtain
$P\!\left(\max_{\alpha\in M_{Q}}\left|\partial^{\alpha}s^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-\partial^{\alpha}f(\hat{x})\right|\geq\mathcal{K}\,h_{n}^{k-m}\log n\right)$
$\qquad\leq\sum_{\alpha\in M_{Q}}P\!\left(\left|\partial^{\alpha}s^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-\partial^{\alpha}f(\hat{x})\right|\geq\mathcal{K}\,h_{n}^{k-|\alpha|}\log n\right)\leq\frac{2}{n^{c}}+\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}+\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$
since $(M_{Q})$ is finite. Consequently,
$P\!\left(\left|Qs^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-Qf(\hat{x})\right|\geq\mathcal{K}\,h_{n}^{k-m}\log n\right)\leq\frac{2}{n^{c}}+\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}+\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}$
i.e.,
$P\!\left(\left|Qs^{\text{MLS}}_{F_{n},X_{n}}(\hat{x})-Qf(\hat{x})\right|\geq\mathcal{K}\,h_{n}^{k-m}\log n\right)\;\xrightarrow[n\to\infty]{}\;0,$
which proves the theorem. ∎
In order to prove Theorem 2, we rely on a series of auxiliary lemmas. These lemmas provide probabilistic bounds for key components of the MLS method, including the weight function, the matrix $\mathscr{A}_{n}(x)$, its inverse, and the shape functions. The results are used to control the behavior of the approximation and establish the required convergence properties.
We begin by controlling the derivatives of the matrix $\mathscr{A}_{n}(x)$, which appears in the normal equations of the MLS method (39).
Remark 5.
For simplicity of the writing, in Lemmas (5), (7), and (8), the point $x\in\Omega$ is assumed to be an interior point of $\Omega$. For points on the boundary, the results remain valid, due to the interior cone condition, although the constants may differ. By taking the smaller value for the lower bound and the larger value for the upper bound, the statements can be extended to cover all cases.
Lemma 5.
Let the event $\mathscr{E}_{I}$ occurs (35). Then, for a fixed but arbitrary evaluation point ${x}\in\Omega$, and for every $\varepsilon>0$ we have
$P(|\partial^{\alpha_{0}}\mathscr{A}_{n}({x})-C_{\alpha}h_{n}^{-|\alpha_{0}|}|\leq\varepsilon\Big|\mathscr{E}_{I}\,\,\forall\,\,x\in\Omega\,\,\text{and}\,\,\,\alpha,\beta\in A)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}},$ (45)
where $C_{\alpha}$ is a positive constant, and $\mathscr{A}_{n}$ is a matrix which its elements are given in (39).
Proof.
Let $B\left(\hat{x},h_{n}\right)=\left\{x\in\mathbb{R}^{d}:\|x-\hat{x}\|\leq h_{n}\right\}$ where $h_{n}$ is the fill distance as defined in Theorem 3 satisfying in probability that $h_{n}\rightarrow 0\text{ as }n\rightarrow\infty$. set $I_{B\left(\hat{x},h_{n}\right)}=\left\{i:X_{i}\in B(\hat{x},h_{n})\right\}$ and $N_{B\left(\hat{x},h_{n}\right)}=\#I_{B\left(\hat{x},h_{n}\right)}$ the number of elements in this ball. For $\Pi_{k-1}^{d}$ we consider the basis
$p_{\alpha}(x)=\left\{\frac{(x-\hat{x})^{\alpha}}{h_{n}^{|{\alpha}|}}\right\}_{0\leqslant|{\alpha}|\leq k-1},$ (46)
which is the monomial bases centered around the point $\hat{x}$ and scaled by the fill distance. This choice of scaled monomial basis is made for convenience. The MLS formulation, and in particular the matrix $\mathscr{A}_{n}$ and its derivatives, are invariant under a change of basis in $\Pi_{k-1}^{d}$. Consequently, the estimates derived below do not depend on the specific choice of polynomial basis and hold for any basis of $\Pi_{k-1}^{d}$. Thus, the MLS function can be written as
$s^{\text{MLS}}_{F_{n},X_{n}}(x)=\sum_{\alpha\in A}\hat{b}_{n\alpha}(x)p_{\alpha}(x),$ (47)
where $A$ denotes the collection of all $d$-tuples $\alpha$ of nonnegative integers such that $|\alpha|\leq k-1$. The expression $\hat{b}_{n\alpha}(x)$ is the $\alpha$-coordinate of the vector of size $A$ which is given by
$\hat{b}_{n}(x)=\mathscr{A}_{n}^{-1}(x)\mathscr{X}_{n}^{\top}(x)F_{n},$ (48)
where $\mathscr{X}_{n}(x)$ and $\mathscr{A}_{n}(x)$ are defined as follows: Using our basis defined in Equation (46), we can construct the normal equations accordingly (The full formulation of the least-squares problem and the solution of the resulting normal equations are given in the Appendix 8.3). The $n\times A$ design matrix, $\tilde{\mathscr{X}}_{n}(x)$, takes the form
$\tilde{\mathscr{X}}_{n}(x)=\left(\begin{array}[]{cccc}\sqrt{\theta_{1}}p_{\alpha_{0}}(X_{1})&\sqrt{\theta_{1}}p_{\alpha_{1}}(X_{1})&\cdots&\sqrt{\theta_{1}}p_{\alpha_{|A|}}(X_{1})\\ \vdots&&&\vdots\\ \sqrt{\theta_{n}}p_{\alpha_{0}}(X_{n})&\sqrt{\theta_{n}}p_{\alpha_{1}}(X_{n})&\cdots&\sqrt{\theta_{n}}p_{\alpha_{|A|}}(X_{n})\end{array}\right),$
where $\theta_{i}=\theta_{h}(x_{i}-x)$ and $p_{\alpha_{j}};\,\,0\leq j\leq A$ is according to (46). From this, the normal equations are obtained as
$\tilde{\mathscr{X}}_{n}(x)^{\top}\tilde{\mathscr{X}}_{n}(x)\hat{b}_{n}(x)=\tilde{\mathscr{X}}_{n}(x)^{\top}F_{n}(x),$
which leads to the solution
$\hat{b}_{n}(x)=(\tilde{\mathscr{X}}_{n}(x)^{\top}\tilde{\mathscr{X}}_{n}(x))^{-1}\tilde{\mathscr{X}}_{n}(x)^{\top}F_{n}(x).$
Next, we normalize the design matrix by rescaling it:
$\mathscr{X}_{n}(x)=\frac{1}{N_{B\left(x,h_{n}\right)}}\tilde{\mathscr{X}}_{n}(x),$
i.e.,
$\mathscr{X}_{ni\alpha}(x)=\begin{cases}\frac{\sqrt{\theta_{h}(x_{i}-x)}}{N_{B\left(x,h_{n}\right)}}\frac{\left(x_{i}-x\right)^{\alpha}}{h_{n}^{|\alpha|}},&i\in I_{B\left(x,h_{n}\right)}\text{ and }\alpha\in A,\\ 0,&i\notin I_{B\left(x,h_{n}\right)}\text{ and }\alpha\in A.\end{cases}$
Using this notation, let $\mathscr{A}_{n}(x)$ be the $A\times A$ matrix defined by
$\mathscr{A}_{n}(x)$ $=\frac{1}{N_{B\left(x,h_{n}\right)}}\cdot\tilde{\mathscr{X}}_{n}(x)^{\top}\tilde{\mathscr{X}}_{n}(x)$
$=\frac{1}{N_{B\left(x,h_{n}\right)}}\cdot N_{B\left(x,h_{n}\right)}^{2}\cdot\mathscr{X}_{n}(x)^{\top}\mathscr{X}_{n}(x)$
$=N_{B\left(x,h_{n}\right)}\cdot\mathscr{X}_{n}(x)^{\top}\mathscr{X}_{n}(x).$
Therefore, each matrix element is given by
$\mathscr{A}_{n\alpha\beta}(x)=\frac{1}{N_{B\left(x,h_{n}\right)}}\sum_{I_{B\left(x,h_{n}\right)}}\theta_{h}(x_{i}-x)\frac{\left(x_{i}-x\right)^{\alpha}\left(x_{i}-x\right)^{\beta}}{h_{n}^{|\alpha|+|\beta|}}.$ (49)
With these notations, we obtain the terms $\mathscr{X}_{n}(x)$, $\mathscr{A}_{n}(x)$ and $\hat{b}_{n}(x)$. Now, note that $\mathscr{A}_{n}$ represents the empirical weighted covariance matrix of the random variable (i.e., the sample points $x_{1},...,x_{n}$) expressed in the polynomial basis. Define the normalized weight function by
$w(x_{i}-x)=\frac{\theta_{h}(x_{i}-x)}{1/h_{n}\int_{B(0,sh_{n})}\theta_{h}(u-x)du}=\frac{\Phi(x_{i}/h_{n}-x/h_{n})}{1/h_{n}\int_{B(0,sh_{n})}\Phi(u/h_{n}-x/h_{n})du},$
using change of variables with $\left[t=\frac{u-x}{h_{n}},u-x=th_{n},du=h_{n}dt\right]$ we obtain,
$\frac{\Phi(t_{i})}{1/h_{n}\int_{t\in B(0,s)}\Phi(t)h_{n}dt}=\frac{\Phi(t_{i})}{\int_{t\in B(0,s)}\Phi(t)dt}.$
where $t_{i}=\frac{x_{i}-x}{h_{n}}$. Therefore,
$w(x_{i}-x)=\frac{\Phi(t_{i})}{\int_{t\in B(0,s)}\Phi(t)dt},$ (50)
so that $w(x)$ is a distribution over the ball $B(0,sh)$. Hence, the normalized empirical matrix is given by
$\mathscr{A}_{n\alpha\beta}^{w}(x)=N_{B\left(x,h_{n}\right)}\sum_{I_{B\left(x,h_{n}\right)}}w(x_{i}-x)\mathscr{X}_{ni\alpha}(x)\mathscr{X}_{ni\beta}(x).$
Hence,
$\mathscr{A}^{w}_{n}=\mathscr{A}_{n}\cdot\frac{1}{\int_{t\in B(0,s)}\Phi(t)dt}.$ (51)
Define the $A\times A$ matrix $\mathscr{A}^{w}$ by
$\mathscr{A}^{w}_{\alpha\beta}={\int_{|z-x|\leq s}(z-x)^{\alpha}(z-x)^{\beta}w(z-x)dz},$
[…]
By matrix Bernstein inequality, for every $\varepsilon>0$,
$P\left(\left|\mathscr{A}_{na\beta}^{w}(x)-\mathscr{A}^{w}_{\alpha\beta}\right|\leq\varepsilon\Big|\mathscr{E}_{I},\,\forall x\in\Omega\text{ and }\alpha,\beta\in A\right)\geq 1-2e^{-c^{\prime}\cdot\varepsilon^{2}\log(n)}=1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}},$
where $c^{\prime}$ is a positive constant. So if we denote $C^{\Phi}=\int_{t\in B(0,s)}\Phi(t)dt$, then by Equation (51)
$P\left(\left|\mathscr{A}_{na\beta}(x)-C^{\Phi}\cdot\mathscr{A}^{w}_{\alpha\beta}\right|\leq\varepsilon\Big|\mathscr{E}_{I},\,\forall x\in\Omega\text{ and }\alpha,\beta\in A\right)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}.$ (52)
Now, we claim that for all $|\alpha|\leq k$, there exists a constant $C_{\alpha}$ such that
$\left|\partial^{\alpha}w(x_{i}-x)\right|\leq C_{\alpha}h_{n}^{-|\alpha|},\quad\forall x\in\Omega.$ (53)
Since $\Phi(x)$ is a compactly supported and $C^{k}$ function, all of its derivatives up to order $k$ are continuous and bounded (see Lemma 6). Combining Equations (52) and (53) together implies that
$P(|\partial^{\alpha_{0}}\mathscr{A}_{n}({x})-C_{\alpha}h_{n}^{-|\alpha_{0}|}|\leq\varepsilon\Big|\mathscr{E}_{I}\,\,\forall\,\,x\in\Omega\,\,\text{and}\,\,\,\alpha,\beta\in A)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}.$ (54)
Next, we bound the derivatives of the weight function used in the MLS formulation. This is necessary for controlling the growth of the entries in $\mathscr{A}_{n}(x)$ and related expressions.
Lemma 6.
For $w(x_{i}-x)$ a weighted function as defined in (50), we have
$\left|\partial^{\alpha}w(x_{i}-x)\right|\leq C_{\alpha}h_{n}^{-|\alpha|},\quad\forall x\in\Omega.$
Proof.
Applying the chain rule to $w(x)$, we differentiate:
$\partial^{\alpha}w(x_{i}-x)=\partial^{\alpha}\left(\frac{\Phi(x_{i}/h_{n}-x/h_{n})}{\int_{t\in B(0,s)}\Phi(x_{i}/h_{n}-u/h_{n})du}\right).$
Since the denominator is independent of $x$, differentiation only affects the numerator. Using the chain rule,
$\partial^{\alpha}\Phi\left(\frac{x}{h_{n}}\right)=h_{n}^{-|\alpha|}\left(\partial^{\alpha}\Phi\right)\left(\frac{x}{h_{n}}\right).$
Thus,
$\partial^{\alpha}w(x_{i}-x)=\frac{h_{n}^{-|\alpha|}\left(\partial^{\alpha}\Phi\right)(x_{i}/h_{n}-x/h_{n})}{\int_{t\in B(0,s)}\Phi(x_{i}/h_{n}-u/h_{n})du}.$
Since $\Phi$ and its derivatives are bounded, there exists a constant $C_{\alpha}^{\prime}$ such that
$\left|\left(\partial^{\alpha}\Phi\right)(x_{i}/h_{n}-x/h_{n})\right|\leq C_{\alpha}^{\prime}.$
Therefore,
$\left|\partial^{\alpha}w(x_{i}-x)\right|\leq\frac{C_{\alpha}^{\prime}h_{n}^{-|\alpha|}}{\int_{t\in B(0,s)}\Phi(x_{i}/h_{n}-u/h_{n})du}.$
The denominator is independent of $x$ and is just a normalization factor ensuring $w(x_{i}-x)$ integrates to 1 . Then if we denote $C_{\alpha}=\frac{C^{\prime}_{\alpha}}{\int_{t\in B(0,s)}\Phi(x_{i}/h_{n}-u/h_{n})du}$ we obtain
$\left|\partial^{\alpha}w(x_{i}-x)\right|\leq C_{\alpha}h_{n}^{-|\alpha|}.$
To ensure the stability of the MLS solution, we need a lower bound on the smallest eigenvalue of $\mathscr{A}_{n}(x)$, which guarantees invertibility with high probability.
Lemma 7.
Given the event $\mathscr{E}_{I}$, there exists a constant $C_{\lambda}>0$, such that for every $0<\varepsilon<C_{\lambda}$
$P\left(\lambda_{\min}\left(\mathscr{A}_{n}(x)\right)\geq C_{\lambda}-\varepsilon\Big|\mathscr{E}_{I},\,\,\forall x\in\Omega\right)\geq 1-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$
where $c_{\mathcal{Q}}$ is a real positive constant and $\lambda_{\text{min }}\left(\mathscr{A}_{n}(x)\right)$ is the smallest eigenvalue of $\mathscr{A}_{n}(x)$. In particular, $\mathscr{A}_{n}(x)$ is positive definite and full rank for all sufficiently small $h_{n}$.
Proof.
We want to prove that $\mathscr{A}_{n}(x)$ is uniformly positive definite, i.e., its smallest eigenvalue is bounded away from zero independently of $n$, given 1, 49 and under Lemma 3.
Let $p=\left(p_{\alpha}\right)_{\alpha\in A}$ be a unit vector in $\mathbb{R}^{A}$, i.e., $\sum_{\alpha}p_{\alpha}^{2}=1$. Define the polynomial:
$P(z)=\sum_{\alpha\in A}p_{\alpha}z^{\alpha}$
Let us define the rescaled coordinates:
[…]
Now define the empirical quadratic form:
$\mathcal{Q}_{n}(p):=p^{\top}\mathscr{A}_{n}(x)p=\frac{1}{N_{B\left(x,h_{n}\right)}}\sum_{i\in I_{B\left(x,h_{n}\right)}}\theta_{h}\left(x_{i}-x\right)\cdot\left(\sum_{\alpha}p_{\alpha}\frac{\left(x_{i}-x\right)^{\alpha}}{h_{n}^{|\alpha|}}\right)^{2}.$
Substituting $t_{i}=\frac{x_{i}-x}{h_{n}}$, and using $\theta_{h}\left(x_{i}-x\right)=\Phi\left(t_{i}\right)$, we get:
$\mathcal{Q}_{n}(p)=\frac{1}{N_{B\left(x,h_{n}\right)}}\sum_{i\in I_{B\left(x,h_{n}\right)}}\Phi\left(t_{i}\right)\cdot\left(\sum_{\alpha}p_{\alpha}t_{i}^{\alpha}\right)^{2}=\frac{1}{N_{B\left(x,h_{n}\right)}}\sum_{i\in I_{B\left(x,h_{n}\right)}}\Phi\left(t_{i}\right)\cdot\tilde{P}\left(t_{i}\right)^{2},$
where $\tilde{P}(t)=\sum_{\alpha}p_{\alpha}t^{\alpha}$. By Lemma 3, the samples $\left\{t_{i}\right\}_{i\in I_{B\left(x,h_{n}\right)}}$ inside the rescaled ball $B(0,s)$ become dense as $n\rightarrow\infty$ with high probability, i.e., for any fixed polynomial $p$, the empirical quadratic form
$\mathcal{Q}_{n}(p)=\frac{1}{N_{B\left(x,h_{n}\right)}}\sum_{i\in I_{B\left(x,h_{n}\right)}}\Phi\left(t_{i}\right)\cdot\tilde{P}\left(t_{i}\right)^{2},$
satisfies with high probability
$\mathcal{Q}_{n}(p)\rightarrow\mathcal{Q}(p):=\int_{B(0,s)}\Phi(t)\cdot\tilde{P}(t)^{2}dt.$
More precisely, by Hoffding’s inequality, for every $\varepsilon>0$,
$P\left(\left|\mathcal{Q}_{n}(p)-\mathcal{Q}(p)\right|\leq\varepsilon\Big|\mathscr{E}_{I}\right)\geq 1-2e^{-c_{\mathcal{Q}}\varepsilon^{2}\log n}=1-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$
where $c_{\mathcal{Q}}$ is a positive real constant. Now we show that $\mathcal{Q}(p)$ is strictly positive. Suppose, for contradiction, that $\inf_{p}\mathcal{Q}(p)=0$. Then there is a sequence of polynomials $P_{m}(t)=$ $\sum_{\alpha}p_{\alpha}^{(m)}t^{\alpha}$ with $\left\|p^{(m)}\right\|=1$, such that $Q\left(P_{m}\right)\rightarrow 0$, i.e.,
$\int_{B(0,s)}\Phi(t)\cdot P_{m}(t)^{2}dt\rightarrow 0.$
Since $\Phi(t)>0$ on $B(0,s)$, this implies $P_{m}\rightarrow P$ in $L^{2}(B(0,s))$, and by compactness (unit sphere is compact in finite-dimensional space), there is a limit point $P\neq 0$ such that
$\int_{B(0,s)}P(t)^{2}dt=0,$
which implies $P\equiv 0$, contradicting $\|p\|=1$. Hence, there exists a constant $C_{\lambda}$ such that
$\mathcal{Q}(p)\geq C_{\lambda}>0\quad\text{ for all unit }p.$
Now, recall that by definition:
$\lambda_{\min}\left(\mathscr{A}_{n}(x)\right)=\inf_{\|p\|=1}\mathcal{Q}_{n}(p).$
Therefore, by combining the deterministic positivity of $\mathcal{Q}(p)$ with the stochastic convergence of $\mathcal{Q}_{n}(p)$ to $\mathcal{Q}_{p}$ we conclude that for every $0<\varepsilon<C_{\lambda}$,
$P\left(\lambda_{\min}\left(\mathscr{A}_{n}(x)\right)\geq C_{\lambda}-\varepsilon\Big|\mathscr{E}_{I},\,\,\forall x\in\Omega\right)\geq 1-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$
which means the minimum eigenvalue of $\mathscr{A}_{n}(x)$ is bounded away from zero with high probability. ∎
Having established the invertibility of $\mathscr{A}_{n}(x)$ and bounds on its derivatives, we now derive corresponding bounds for the derivatives of its inverse. This result relies on the two previous Lemmas 5 and 7.
Lemma 8.
Given the event $\mathscr{E}_{I}$ hold, and under the assumptions of Lemma 5 and Lemma 7, then for all $0<\varepsilon<C_{\lambda}$, and $\forall x\in\Omega\,\,\text{and}\,\,\,\alpha,\beta\in A,$
$P\left(\left|\partial^{\alpha_{0}}\mathscr{A}_{n}^{-1}(x)\right|\leq C_{\alpha_{0}}h_{n}^{-\left|\alpha_{0}\right|}\Big|\mathscr{E}_{I}\right)\geq 1-\frac{2}{n^{c^{\prime}\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$ (55)
where $C_{\alpha_{0}}$ is a positive constant depends on $\varepsilon$ but independent of $h_{n}$.
Proof.
We consider the matrix identity
$\mathscr{A}_{n}^{-1}(x)\mathscr{A}_{n}(x)=I.$
Differentiating both sides with respect to a multi-index $\alpha_{0}$ and evaluating at a point $x\in\Omega$, we obtain
$\partial^{\alpha_{0}}\mathscr{A}_{n}^{-1}(x)=-\sum_{\beta<\alpha_{0}}\binom{\alpha_{0}}{\beta}\mathscr{A}_{n}^{-1}(x)\left(\partial^{\alpha_{0}-\beta}\mathscr{A}_{n}(x)\right)\left(\partial^{\beta}\mathscr{A}_{n}^{-1}(x)\right).$
We now prove the desire result by induction on $\left|\alpha_{0}\right|$.
Starting with the base case: $\left(\left|\alpha_{0}\right|=1\right)$. Let $\alpha_{0}=e_{j}$, the unit multi-index in the $j$ th direction. Then,
$\partial^{e_{j}}\mathscr{A}_{n}^{-1}(x)=-\mathscr{A}_{n}^{-1}(x)\left(\partial^{e_{j}}\mathscr{A}_{n}(x)\right)\mathscr{A}_{n}^{-1}(x).$
From Lemma 5 we have,
$P\left(|\partial^{e_{j}}\mathscr{A}_{n}(x)-C_{\alpha}h_{n}^{-1}|\leq\varepsilon\Big|\mathscr{E}_{I}\right)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}},$
and from Lemma 7, we know that given the event $\mathscr{E}_{I}$, the matrix $\mathscr{A}_{n}(x)$ is a positive definite with probability at least $1-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}$, hence:
$P\left(\left\|\mathscr{A}_{n}^{-1}(x)\right\|_{2}\leq\frac{1}{C_{\lambda}-\varepsilon}\Big|\mathscr{E}_{I}\right)\geq 1-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}.$ (56)
Using the standard bound $|M|\leq C\|M\|_{2}$ for matrices, we conclude:
$P\left(|\partial^{e_{j}}\mathscr{A}_{n}^{-1}(x)|\leq C_{e_{j}}h_{n}^{-1}\Big|\mathscr{E}_{I}\right)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$
where $C_{e_{j}}h_{n}^{-1}=\frac{C_{\alpha}h_{n}^{-1}+\varepsilon}{(C_{\lambda}-\varepsilon)^{2}}$. This establishes the base case. Next, the inductive step is assume that for all multi-indices $\beta<\alpha_{0}$, we have for all $x\in\Omega$
$P\left(|\partial^{\beta}\mathscr{A}_{n}^{-1}(x)|\leq C_{\beta}h_{n}^{-|\beta|}\Big|\mathscr{E}_{I}\right)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}.$ (57)
Now we prove the inductive step. From the derivative identity we obtain:
$\left|\partial^{\alpha_{0}}\mathscr{A}_{n}^{-1}(x)\right|\leq\sum_{\beta<\alpha_{0}}\binom{\alpha_{0}}{\beta}\left\|\mathscr{A}_{n}^{-1}(x)\right\|_{2}\cdot\left|\partial^{\alpha_{0}-\beta}\mathscr{A}_{n}(x)\right|\cdot\left|\partial^{\beta}\mathscr{A}_{n}^{-1}(x)\right|.$ (58)
Using the induction hypothesis (57) together with Equations (56),(58), we obtain that,
$P\left(|\partial^{\alpha_{0}}\mathscr{A}_{n}^{-1}({x})|\leq C_{\alpha_{0}}h_{n}^{-|\alpha_{0}|}\Big|\mathscr{E}_{I}\right)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$
where $C_{\alpha_{0}}(\varepsilon)\sim\frac{1}{\left(C_{\lambda}-\varepsilon\right)^{k+1}}$. This completes the inductive argument, as desired.
Finally, we use the preceding bounds to control the derivatives of the shape functions $a_{i}^{*}(x)$ themselves. This establishes a uniform stability estimate essential for proving convergence in the MLS approximation. The following Lemma is the stochastic version of Theorem 3.11 in Mirzaei (2015).
Lemma 9.
Given that the event $\mathscr{E}_{I}$ holds, then for all $0<\varepsilon<C_{\lambda}$ the MLS shape functions $a^{*}_{i}$ of (38) have the following stability condition for $|\alpha|\leq k-1$ and for all $x\in\Omega$,
$P\left(\sum_{i\in I_{B\left(x,sh_{n}\right)}}\left|\partial^{\alpha}a_{i}^{*}(x)\right|\leq C_{\alpha}^{\prime}\cdot\log(n)\cdot h_{n}^{-|\alpha|}\Big|\mathscr{E}_{I}\right)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$ (59)
where $C^{\prime}_{\alpha}\sim\frac{1}{\left(C_{\lambda}-\varepsilon\right)^{|\alpha|+1}}$ is independent of $x$, and provided that all assumptions of Lemma 6 are satisfied.
Proof.
Recall the stochastic MLS shape functions (38) are given by
$a_{i}^{*}(x)=\theta_{h}(x_{i}-x)\sum_{|\beta|\leq k-1}\eta_{\beta}(x)\cdot\frac{\left(x_{i}-x\right)^{\beta}}{h_{n}^{|\beta|}}.$ (60)
We start with bounding the derivatives of $\eta(x)=\mathscr{A}_{n}(x)^{-1}\boldsymbol{p}(x)$.
$\partial^{\alpha}\eta(x)=\sum_{\zeta\leq\alpha}\binom{\alpha}{\zeta}\partial^{\alpha-\zeta}\mathscr{A}_{n}^{-1}(x)\cdot\partial^{\zeta}\boldsymbol{p}(x)\quad\forall\alpha\text{ with }|\alpha|\leqslant k-1.$
Now, by the definition of the scaled polynomial basis, we have:
$\partial^{\zeta}\boldsymbol{p}(x)=\zeta!\cdot h_{n}^{-|\zeta|}\cdot e_{\zeta},$
where $e_{\zeta}$ is the canonical basis vector in multi-index ordering (i.e., the standard basis vector with a 1 in the position corresponding to multi-index $\zeta$, and zeros elsewhere). By Lemma 8, with probability of at least $1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}$, the inverse matrix satisfies
$\left|\partial^{\gamma}\mathscr{A}_{n}^{-1}(x)\right|\leq C_{\gamma}h_{n}^{-|\gamma|},$
where $C^{\prime}_{\gamma}\sim\frac{1}{\left(C_{\lambda}-\varepsilon\right)^{|\gamma|+1}}$ Hence, with the same probability, for a constant matrix $C_{\alpha,\zeta}$:
$\left|\partial^{\alpha}\eta(x)\right|\leq\sum_{\zeta\leq\alpha}\binom{\alpha}{\zeta}{C_{\alpha,\zeta}h_{n}^{-|\alpha-\zeta|}}\cdot e_{\zeta}h_{n}^{-|\zeta|}\leq C_{\alpha}h_{n}^{-|\alpha|},$ (61)
where the vector $C_{\alpha}$ is a bound for $\sum_{\zeta\leqslant\alpha}\binom{\alpha}{\zeta}C_{\alpha,\zeta}e_{\zeta}$. Taking the derivatives of both sides in Equation 60 and using the product rule, we have,
$\partial^{\alpha}a_{i}^{*}(x)=\sum_{\zeta\leq\alpha}\binom{\alpha}{\zeta}\partial^{\alpha-\zeta}\theta_{h}(x_{i}-x)\sum_{|\beta|\leq k-1}\partial^{\zeta}\eta_{\beta}(x)\cdot\frac{\left(x_{i}-x\right)^{\beta}}{h_{n}^{|\beta|}}.$
By Lemma 6, the weight functions satisfy:
$\left|\partial^{\alpha}\theta_{h}(x_{i}-x)\right|\leq C_{\alpha,\theta}h_{n}^{-|\alpha|},$
and since $x_{i}\in B\left(x,sh_{n}\right)$, we have $\left|\left(x_{i}-x\right)^{\beta}\right|\leq\left(sh_{n}\right)^{|\beta|}$. Applying the latter in addition to the previous bounds, we get that with probability of at least $1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}}$,
$\left|\partial^{\alpha}a_{i}^{*}(x)\right|\leqslant\sum_{\zeta\leqslant\alpha}\left\{\binom{\alpha}{\zeta}C_{\alpha,\zeta}h_{n}^{|\zeta|-|\alpha|}\sum_{|\beta|\leqslant k-1}C_{\alpha}h_{n}^{-|\zeta|}h_{n}^{-|\beta|}h_{n}^{|\beta|}\right\}\leqslant C_{\alpha}h_{n}^{-|\alpha|}.$
Apply Lemma 3 yields that for all $x\in\Omega$ and for all $0<\varepsilon<C_{\lambda}$,
$P\left(\sum_{i\in I_{B\left(x,sh_{n}\right)}}\left|\partial^{\alpha}a_{i}^{*}(x)\right|\leq C_{\alpha}^{\prime}\cdot\log(n)\cdot h_{n}^{-|\alpha|}\Big|\mathscr{E}_{I}\right)\geq 1-\frac{2}{n^{c^{\prime}\cdot\varepsilon^{2}}}-\frac{2}{n^{c_{\mathcal{Q}}\varepsilon^{2}}},$ (62)
where $C^{\prime}_{\alpha}\sim\frac{1}{\left(C_{\lambda}-\varepsilon\right)^{|\alpha|+1}}$, as desired. ∎
6 Smoothness of the Approximation
In this section, we show that the MLS approximation $s^{\text{MLS}}_{F_{n},X_{n}}(x)$ inherits smoothness from the weight function and the local sample configuration. The structural assumptions on the point cloud ensure that, with high probability, each local neighborhood contains sufficiently many well-distributed points. This guarantees that the local least-squares matrix $\mathscr{A}_{n}(x)$ is full-rank and smoothly invertible. Together with the weight function $\theta_{h}\in\mathcal{C}^{k}$, this implies that the shape functions $a_{i}^{*}(x)$ inherit the same smoothness. Consequently, the MLS approximation $s^{\text{MLS}}_{F_{n},X_{n}}(x)$, expressed as a linear combination of the sampled values $f(x_{i})$ with these smooth weights, belongs to $\mathcal{C}^{k}(\Omega)$ once the theorem is applied.
Proof of Theorem 3.
Let $\theta_{h}\in\mathcal{C}^{k}$. We show that, with probability one as $n\to\infty$, the MLS approximation $s^{\text{MLS}}_{F_{n},X_{n}}(x)$ belongs to $\mathcal{C}^{k}(\Omega)$.
Set $\varepsilon=\frac{C_{\lambda}}{10}$, then by Lemma 7, for every fixed $x_{0}\in\Omega$,
$P\!\big(\lambda_{\min}(\mathscr{A}_{n}(x_{0}))\geq\frac{9C_{\lambda}}{10}\big)\geq 1-\frac{2}{n^{\tilde{c}_{\lambda}}},$
where $\tilde{c}_{\lambda}=c_{\mathcal{Q}}\cdot(\frac{C_{\lambda}}{10})^{2}$. To show that $s^{\text{MLS}}_{F_{n},X_{n}}(x)$ is $\mathcal{C}^{k}$ on all of $\Omega$, we need the matrix $\mathscr{A}_{n}(x)$ to be uniformly invertible (positive definite) and smooth across the domain. Lemma 7 only gives a pointwise probabilistic lower bound on $\lambda_{\min}(\mathscr{A}_{n}(x_{0}))$. To extend this to all $x\in\Omega$, we apply Weyl’s eigenvalue perturbation inequality for Hermitian matrices, which allows us to control how the minimal eigenvalue changes with $x$ as follows
$|\lambda_{\min}(\mathscr{A}_{n}(x_{0}))-\lambda_{\min}(\mathscr{A}_{n}(x))|\leq\|\mathscr{A}_{n}(x_{0})-\mathscr{A}_{n}(x)\|_{op}\quad$
Since all matrix norms on $\mathbb{R}^{d\times d}$ are equivalent, there exists a constant $C_{\text{norm}}>0$ such that
$\|B\|_{op}\leq C_{\text{norm}}\|B\|_{F,1},$
where $\|B\|_{F,1}=\sum_{i,j}|B_{ij}|$ is the entrywise Frobenius 1-norm. Recall (39) then $\mathscr{A}_{n,ij}(x)$ is given by,
$\mathscr{A}_{n,ij}(x)=\sum_{t=1}^{n}\theta_{h}\left(x_{t}-x\right)\frac{\left(x_{t}-x\right)^{\alpha_{i}}}{h_{n}^{\left|\alpha_{i}\right|}}\frac{\left(x_{t}-x\right)^{\alpha_{j}}}{h_{n}^{\left|\alpha_{j}\right|}},\quad i,j=1,\ldots,A.$
Each entry $\mathscr{A}_{n,ij}(x)$ of $\mathscr{A}_{n}(x)$ is $\mathcal{C}^{k}$ as a function of $x$, since it is built from smooth weights $\theta_{h}$ and polynomial basis functions. Therefore, each $\mathscr{A}_{n,ij}(x)$ is locally Lipschitz, and on the compact domain $\Omega$ we can choose a uniform Lipschitz constant $C_{\text{Lip}}$ such that
$\sum_{i,j}|\mathscr{A}_{n,ij}(x_{0})-\mathscr{A}_{n,ij}(x)|\leq C_{\text{Lip}}\cdot d\cdot\|x_{0}-x\|.$
Combining the above inequalities, we obtain a Lipschitz bound for the minimal eigenvalue:
$|\lambda_{\min}(\mathscr{A}_{n}(x_{0}))-\lambda_{\min}(\mathscr{A}_{n}(x))|\leq C_{\text{norm}}\cdot C_{\text{Lip}}\cdot d\cdot\|x_{0}-x\|.$
Let
$\delta_{0}=\frac{C_{\lambda}}{2}\cdot\frac{1}{C_{\text{norm}}\cdot C_{\text{Lip}}\cdot d}.$
Then, for all $x\in B(x_{0},\delta_{0})$, assuming that the event $\left\{\lambda_{\min}\left(\mathscr{A}_{n}\left(x_{0}\right)\right)\geq\frac{9C_{\lambda}}{10}\right\}$ holds,
$|\lambda_{\min}(\mathscr{A}_{n}(x_{0}))-\lambda_{\min}(\mathscr{A}_{n}(x))|<\frac{C_{\lambda}}{2}\quad\implies\quad\lambda_{\min}(\mathscr{A}_{n}(x))>\frac{C_{\lambda}}{2}.$
Hence, for each fixed $x_{0}$,
$P\!\Big(\lambda_{\min}(\mathscr{A}_{n}(x))\geq\tfrac{C_{\lambda}}{2}\text{ for all }x\in B(x_{0},\delta_{0})\Big)\geq 1-\frac{2}{n^{\tilde{c}_{\lambda}}}.$
Now, since $\Omega$ is compact, it can be covered by finitely many balls $B(y_{j},\delta_{0})$, $j=1,\dots,M$ where $\delta_{0}>0$ is independent of $n$, $y_{i}$ are arbitrary points in $\Omega$ and $M$ is the minimal number of balls of radius $\delta_{0}$ required to cover $\Omega$ (see (9)). We consider, for each $j=1,\ldots,M$, the events
$E_{j}=\left\{\lambda_{\min}\left(\mathscr{A}_{n}(x)\right)\geq\frac{C_{\lambda}}{2}\text{ for all }x\in B\left(y_{j},\delta_{0}\right)\right\}.$
Since each of these events $E_{j}$ occurs with probability at least $1-\frac{2}{n^{\tilde{c}_{\lambda}}}$, the union bound yields
$P\left(E_{1}\cap\cdots\cap E_{M}\right)\geq 1-M\cdot\frac{2}{n^{\tilde{c}_{\lambda}}},$
i.e.,
$P\!\Big(\forall j=1,\dots,M:\lambda_{\min}(\mathscr{A}_{n}(x))\geq\tfrac{C_{\lambda}}{2}\text{ for all }x\in B(y_{j},\delta_{0})\Big)\geq 1-M\cdot\frac{2}{n^{\tilde{c}_{\lambda}}}.$
Therefore, with probability tending to one, the matrix $\mathscr{A}_{n}(x)$ is uniformly positive definite on $\Omega$. Now, recall that the shape functions are defined as (38)
$a_{i}^{*}(x)=\theta_{h}(x_{i}-x)\sum_{|\alpha|\leq k-1}\eta_{\alpha}(x)\frac{(x_{i}-x)^{\alpha}}{h_{n}^{|\alpha|}}.$
Since the entries of $\mathscr{A}_{n}(x)$ are $\mathcal{C}^{k}$, the entries of $\mathscr{A}_{n}(x)^{-1}$ are also $\mathcal{C}^{k}$, and the coefficients $\eta_{\alpha}(x)$, which are the components of $\eta(x)=\mathscr{A}_{n}(x)^{-1}\boldsymbol{p}(x)$ inherit this smoothness. Combining this with the smoothness of the weight $\theta_{h}$, each shape function $a_{i}^{*}(x)$ is $\mathcal{C}^{k}$, and therefore the MLS approximation $s^{\text{MLS}}_{F_{n},X_{n}}(x)\in\mathcal{C}^{k}(\Omega)$ with probability one as $n\to\infty$. ∎
7 Manifold Estimation
Section 2 addresses the problem of function estimation. A natural question that arises is whether this framework can be extended to the setting of manifolds. In this section, we review the work of Sober and Levin Sober and Levin (2020), who consider deterministic sampling assumptions consistent with the function estimation works mentioned above. Their results show that approximation error rate matches the rate obtained in the function approximation case, as established in Levin (1998). We summarize the key components of their analysis and show that, in a similar manner, replacing the deterministic sampling assumptions with stochastic ones preserves the same convergence rate as in Equation 37.
The manifold MLS (MMLS) computation as presented in Sober and Levin (2020), involves two key steps. First, we establish a local coordinate system over which the manifold is locally represented as the graph of a function. Second, we compute a weighted least-squares estimate to approximate the manifold as a function over this local coordinate system. In details, assume that $\mathcal{M}\in\mathcal{C}^{k}(k\geq 2)$ is a closed (i.e., compact and boundary-less) submanifold of $\mathbb{R}^{D}$, where $\mathcal{C}^{k}$ means that the manifold is smooth, with a structure that allows for $k$ continuous derivatives in the usual topology, giving it a well-defined notion of local smoothness. Let $R=\left\{r_{i}\right\}_{i=1}^{n}\subset\mathcal{M}$ be a sample set with respect to the domain $\mathcal{M}$. Then, Given a point $r$ near $\mathcal{M}$ the MMLS projection of $r$ is defined through two sequential steps, following a projective approximation approach, where points in the neighborhood of the manifold are considered and projected onto a nearby local manifold, providing the projection of $r$.
  1. 1.
    Find a local $d$-dimensional affine subspace $H(r)$ centered at a point $q(r)\in\mathbb{R}^{D}$, which serves as an approximation to the sampled data near $r$. Specifically, let $H=q+\operatorname{Span}\left\{e_{k}\right\}_{k=1}^{d}$, where $\left\{e_{k}\right\}_{k=1}^{d}$ is an orthonormal set in $\mathbb{R}^{D}$. This affine subspace will act as a local coordinate system. In other words, we wish to Find a $d$-dimensional affine space $H(r)$, and a point $q(r)$ on $H(r)$, such that
    $(q(r),H(r))=\underset{q\in\mathbb{R}^{D},(H-q)\in Gr(d,D)}{\operatorname{argmin}}J_{1}(q,H\mid r),$ (63)
    where
    $J_{1}(q,H\mid r)=\sum_{i=1}^{n}d\left(r_{i}-q,H\right)^{2}\theta_{1}\left(\left\|r_{i}-q\right\|\right),$ (64)
    under the constraints
    1. (a)
      $r-q\perp H$
    2. (b)
      $q\in B_{\mu}(r)$
    3. (c)
      $\left(R\cap B_{h}(q)\right)\neq 0,$
    where $\theta_{1}(t)$ is a fast decaying radial weight function, $Gr(d,D)$ is the Grassmannian, i.e., the set of $d$-dimensional sub-spaces in an $D$-dimensional vector space, $d\left(r_{i},H\right)$ is the Euclidean distance between the point $r_{i}$ and the affine sub-space $H,\mu$ is some fixed number, $B_{\eta}(x)$ is an open ball of radius $\eta$ around $x$, and $h$ is the fill distance.
  2. 2.
    Define the projection of $r$ using a local polynomial approximant $\pi:H\simeq\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ of $\mathcal{M}$ over the new coordinate system. Explicitly, we define $x_{i}=P_{H}(r_{i}-q)$ as the orthogonal projection of $r_{i}-q$ onto $H\in Gr(d,D)$, and denote the corresponding samples of a function $\varphi$ by $\varphi(x_{i})=r_{i}$. The approximation in the local coordinates is performed by finding a polynomial $\pi^{*}(y)\in\Pi_{k-1}^{d}$ which minimizes
    $J_{2}(p(x)\mid q,H)=\sum_{i=1}^{n}\|p(x_{i})-r_{i}\|^{2}\,\theta_{2}(x_{i}),$ (65)
    where $\theta_{2}(t)$ is a fast-decaying radial weight function consistent across scales, i.e., it is normalized to maintain its relative decay profile across different local scales (e.g., bandwidths). That is,
    $p^{*}_{r}(x\mid r)=\underset{\pi\in\Pi_{k-1}^{d}}{\operatorname{argmin}}J_{2}(\pi(x)\mid q(r),H(r)).$ (66)
    The Manifold-MLS projection is defined as
    $\mathcal{P}_{k}^{h}(r)=p^{*}_{r}(0\mid r)$ (67)
    Using this procedure, the approximating manifold is defined as
    $\mathcal{M}^{h}=\left\{\mathcal{P}_{k}^{h}(p)\mid\forall p\in\mathcal{M}\right\}.$ (68)
Lastly, it is shown that when provided with a finite set of samples of $\mathcal{M}$ with some fill distance h, the MMLS method yields an approximation $\mathcal{M}^{h}$ that converges to $\mathcal{M}$ with a provable rate.
We now state and prove this result in the context of stochastic sampling, employing the notations $h_{n}$ and $\mathcal{M}^{h_{n}}$.
Theorem (Theorem 4).
Assume that $\mathcal{M}^{h_{n}}\in\mathcal{C}^{k}$. Then, there exists a constant $s>0$, independent of $h_{n}$, such that the MMLS approximation using a radial weight function $\theta$ with finite support of size $sh_{n}$ satisfies, for sufficiently small $h_{n}$, the error bound
$\left\|\mathcal{M}^{h_{n}}-\mathcal{M}\right\|_{\text{Hausdorff}}=O_{P}(\log nh_{n}^{k}),$
where $\mathcal{M}^{h_{n}}$ denotes the MMLS approximation of degree k1k-1.
Here, the Hausdorff distance between two sets $\mathcal{A},\mathcal{B}\subset\mathbb{R}^{D}$ is defined as
$\|\mathcal{A}-\mathcal{B}\|_{\text{Hausdorff}}=\max\left\{\sup_{a\in\mathcal{A}}d(a,\mathcal{B}),\sup_{b\in\mathcal{B}}d(b,\mathcal{A})\right\},$
where $d(p,\mathcal{N})$ denotes the Euclidean distance from the point $p$ to the set $\mathcal{N}$.
Proof.
Let $r\in\mathcal{M}^{h_{n}}$. Denote by $(q^{*}(r),H^{*}(r))$ the minimizers of the cost function in equation 63 (step 1 in the algorithm), where $H^{*}(r)$ represents the affine subspace, i.e., the local approximation of the tangent space, and $q^{*}(r)$ represents the origin of the local coordinate system used for the approximation. Now since $H^{*}(r)$ is a first-order (tangent) approximation of the manifold, the deviations from $H^{*}(r)$ scale quadratically with the neighborhood size, i.e.
$d(r_{i},H^{*}(r))=O(h_{n}^{2}).$ (69)
This result yields that $H^{*}(r)$ approximates the sample set $\left\{r_{i}\right\}$ up to the order of $O\left(h_{n}^{2}\right)$ in an $O(h_{n})$ neighborhood of $q^{*}(r)$. Consequently, the objective $J_{1}$ evaluated at the tangent space is an infinite sum of second-order terms:
$J_{1}(q,H\mid r)\geq J_{1}(q^{*}(r),H^{*}(r)\mid r)=\sum_{i=1}^{n}d\left(r_{i}-q^{*}(r),H^{*}(r)\right)^{2}\theta_{1}\left(\left\|r_{i}-q^{*}(r)\right\|\right).$
Letting
$T_{i,n}:=d(r_{i}-q^{*},H^{*})^{2}\,\theta_{1}(\|r_{i}-q^{*}\|),\quad i=1,\dots,n,$
we observe from Equation 69 that each term is $O(h_{n}^{4})$, so that $\sup_{i}T_{i,n}=O(h_{n}^{4})$. Denoting by
$M_{n}:=\#\{i:T_{i,n}\text{ contributes to the sum}\},$
the number of non-negligible terms, then Lemma 3 implies $M_{n}\sim\log n$ with probability at least $1-\frac{2}{n^{c}}$. It then follows that
$\sum_{i=1}^{n}T_{i,n}\leq M_{n}\cdot\sup_{i}T_{i,n}=O_{P}(\log n)\cdot O(h_{n}^{4})=O_{P}(\log n\,h_{n}^{4}),$
and consequently,
$J_{1}(q,H\mid r)\geq J_{1}(q^{*}(r),H^{*}(r)\mid r)=O_{P}(\log n\,h_{n}^{4}).$
Now, according to Theorem 2, as the projection $\mathcal{P}^{h}_{k}\in\mathcal{M}^{h_{n}}$ is merely a local polynomial approximation of $\mathcal{M}$, we achieve that $\mathcal{P}^{h}_{k}$ is $O_{P}(\log nh_{n}^{k})$ away from the manifold $\mathcal{M}$. Accordingly, for all $r\in\mathcal{M},d\left(r,\mathcal{M}^{h_{n}}\right)\leq$ $O_{P}\left(\log nh_{n}^{k}\right)$. Furthermore, for each $s\in{\mathcal{M}}^{h_{n}}$ there exists a point $r\in\mathcal{M}$ such that $s=$ $\mathcal{P}_{k}^{h}(r)$ which is $O_{P}\left(\log nh_{n}^{k}\right)$ away from $\mathcal{M}$. Thus, for all $s\in\mathcal{M}^{h_{n}},d(s,\mathcal{M})\leq O_{P}\left(\log nh_{n}^{k}\right)$ as well, and the theorem follows.
8 Appendix
8.1 Stochastic Boundedness
Definition 5.
The notation
$X_{n}=O_{p}\left(a_{n}\right)\text{ as }n\rightarrow\infty$
means that $\frac{X_{n}}{a_{n}}$ is scholastically bounded. That is, for any ${\varepsilon_{x}}>0$, there exists a finite $M>0$ and a finite $N_{x}>0$ such that
$P\left(\left|\frac{X_{n}}{a_{n}}\right|>M\right)<{\varepsilon_{x}},\,\,\forall n>N_{x}.$
Definition 6.
The notation
$Y_{n}=\Omega_{p}\left(b_{n}\right)\text{ as }n\rightarrow\infty$
means that $\frac{Y_{n}}{b_{n}}$ is scholastically bounded. That is, for any ${\varepsilon_{y}}>0$, there exists a finite $m>0$ and a finite $N_{y}>0$ such that
$P\left(\left|\frac{Y_{n}}{b_{n}}\right|<m\right)<{\varepsilon_{y}},\,\,\forall n>N_{y}.$
8.2 Failure of the deterministic proof in stochastic setup
In Wendland’s work Wendland (2005) (page 42), there is a review of Levin’s work, which proves that $\|s^{\text{MLS}}_{f,X}(x)-f\|_{\Omega,\infty}<L\cdot h^{k}$. The technique used there involves bounding the error by controlling the shape functions through certain manipulations. In order to achieve this, it is necessary to bound the number of samples in $I(x)$ by a constant. This is done by comparing the volume of the union of the small balls with the volume of the large ball, leading to the following bound:
$I(x)\operatorname{vol}(B(0,1))\delta_{n}^{d}\leq\operatorname{vol}(B(0,1))(\delta_{n}+h_{n})^{d}$ (70)
Using the quasi-uniformity property (which holds in the deterministic setup) finally leads to
$I(x)\leq\left(1+\frac{h_{n}}{\delta_{n}}\right)^{d}\leq\left(1+c_{qu}\right)^{d}.$ (71)
This implies that the number of indices in $I(x)$ remains bounded as $n$ tends to infinity, which contradicts Lemma 3, as is the case in stochastic settings. Mirzaei used this result in his paper to prove his theory. He relies on it in several of his lemmas, and therefore his Theorem 3.12—which is the main result we address in our paper (Theorem 2)—cannot be reproduced in the stochastic setting. We re-proved the claims that depend on this result without relying on the quasi-uniformity property, and together with the remaining claims that do not depend on it, we constructed the entire proof.
8.3 Detailed computation of the weighted covariance matrix in proof of Theorem 2
Recall our basis is defined as:
$p_{\alpha}=\left\{\left(\frac{x-z}{h_{n}}\right)^{\alpha}\right\}_{\alpha\in A},$
Using this basis, we construct the normal equations accordingly. The design matrix takes the form
$\tilde{\mathscr{X}}=\left(\begin{array}[]{cccc}\sqrt{\theta_{1}}p_{\alpha_{0}}(X_{1})&\sqrt{\theta_{1}}p_{\alpha_{1}}(X_{1})&\cdots&\sqrt{\theta_{1}}p_{\alpha_{|A|}}(X_{1})\\ \vdots&&&\vdots\\ \sqrt{\theta_{n}}p_{\alpha_{0}}(X_{n})&\sqrt{\theta_{n}}p_{\alpha_{1}}(X_{n})&\cdots&\sqrt{\theta_{n}}p_{\alpha_{|A|}}(X_{n})\end{array}\right),$
where $\theta_{i}=\theta_{h}(xi-x)$. From this, the normal equations are obtained as
$C\cdot\tilde{\mathscr{X}}^{\top}\tilde{\mathscr{X}}\beta=C\cdot\tilde{\mathscr{X}}^{\top}F,$
which leads to the solution
$\beta=(\tilde{\mathscr{X}}^{\top}\tilde{\mathscr{X}})^{-1}\tilde{\mathscr{X}}^{\top}F.$
To explicitly compute $\tilde{\mathscr{X}}^{\top}\tilde{\mathscr{X}}$, we expand the expression:
$\tilde{\mathscr{X}}^{\top}\tilde{\mathscr{X}}=\left(\begin{array}[]{ccc}\sqrt{\theta_{1}}p_{\alpha_{0}}(X_{1})&\cdots&\sqrt{\theta_{n}}p_{\alpha_{0}}(X_{n})\\ \vdots&&\vdots\\ \sqrt{\theta_{1}}p_{\alpha_{|A|}}(X_{1})&\cdots&\sqrt{\theta_{n}}p_{\alpha_{|A|}}(X_{n})\end{array}\right)\cdot\left(\begin{array}[]{ccc}\sqrt{\theta_{1}}p_{\alpha_{0}}(X_{1})&\cdots&\sqrt{\theta_{1}}p_{\alpha_{|A|}}(X_{1})\\ \vdots&&\vdots\\ \sqrt{\theta_{n}}p_{\alpha_{0}}(X_{n})&\cdots&\sqrt{\theta_{n}}p_{\alpha_{|A|}}(X_{n})\end{array}\right)$
$\left(\begin{array}[]{ccc}\sum_{i}{\theta_{i}}p_{\alpha_{0}}(X_{i})\cdot p_{\alpha_{0}}(X_{i})&\cdots&\sum_{i}{\theta_{i}}p_{\alpha_{0}}(X_{i})\cdot p_{\alpha_{|A|}}(X_{i})\\ \vdots&&\vdots\\ \sum_{i}{\theta_{i}}p_{\alpha_{|A|}}(X_{i})\cdot p_{\alpha_{0}}(X_{i})&\cdots&\sum_{i}{\theta_{i}}p_{\alpha_{|A|}}(X_{i})p_{|A|}(X_{i})\end{array}\right).$
Next, we introduce the normalization by defining $N=N_{B\left(x,h_{n}\right)}$ and rescaling the design matrix:
$\mathscr{X}=\frac{1}{N}\tilde{\mathscr{X}}$,
Using this notation, we express $A$ as
$A=\frac{1}{N}\cdot\tilde{\mathscr{X}}^{\top}\tilde{\mathscr{X}}=\frac{1}{N}\cdot N^{2}\cdot\mathscr{X}^{\top}\mathscr{X}=N\cdot\mathscr{X}^{\top}\mathscr{X}.$
With these definitions, we recover the terms $\mathscr{X}_{n}(x)$ and $\mathscr{A}_{n}(x)$ introduced in section 3. Finally, regarding the term $Qs^{\text{MLS}}_{F_{n},X_{n}}$, we note that it is a scalar quantity. To see this, we examine its components:
  1. 1.
    $\mathscr{Q}_{n}^{\top}$ :
    $\mathscr{Q}_{n}^{\top}=\left[\frac{\alpha_{1}!q_{\alpha_{1}}}{h_{n}^{\left[\alpha_{1}\right]}},\frac{\alpha_{2}!q_{\alpha_{2}}}{h_{n}^{\left[\alpha_{2}\right]}},\ldots,\frac{\alpha_{\#(\mathscr{N})}!q_{\alpha_{\#(\mathscr{N})}}}{h_{n}^{\left[\alpha_{\#(\mathscr{N})}\right]}}\right].$
  2. 2.
    $\mathscr{A}_{n}^{-1}(x)$ : General elements of $\mathscr{A}_{n}^{-1}(x)$ are denoted by $\mathscr{A}_{n\alpha\beta}^{-1}(x)$, have the form of $\tilde{\mathscr{X}}^{\top}\tilde{\mathscr{X}}$ up to multiplication by a general constant.
  3. 3.
    $\mathscr{X}_{n}^{\top}(x)$ : The transpose of $\mathscr{X}_{n}(x)$, where:
    $\mathscr{X}_{ni\alpha}(x)=\begin{cases}\frac{\sqrt{\theta_{h}(x_{i}-x)}}{N_{B\left(x,h_{n}\right)}}\frac{\left(x_{i}-x\right)^{\alpha}}{h_{n}^{|a|}},&i\in I_{B\left(x,h_{n}\right)}\text{ and }\alpha\in\mathscr{N},\\ 0,&i\notin I_{B\left(x,h_{n}\right)}\text{ and }\alpha\in\mathscr{N}.\end{cases}$
By multiplying these three terms, we obtain
$\sum_{\alpha\in A}\sum_{\beta\in A}\sum_{i\in I_{B\left(x,h_{n}\right)}}\frac{1}{N_{B\left(x,h_{n}\right)}}\cdot\left[\mathscr{A}_{n}^{-1}(x)\right]_{\alpha\beta}\cdot\frac{\left(x_{i}-x\right)^{\beta}}{h_{n}^{|\beta|}}\cdot\frac{\alpha!q_{\alpha}}{h_{n}^{|\alpha|}}\cdot\left(F_{n}(x)\right).$ (72)
9 Acknowledgments
This work was made possible through the support of the Ariane de Rothschild Women Doctoral Program for outstanding female PhD students.